### v1

# On the multiplier spectrum of polynomials

Abstract.
We prove several results on multiplier spectrum for polynomials. We provide a detailed proof of the theorem stating that the multiplier spectrum morphism is generically injective on the moduli space of polynomials. We obtain a description of the non-injective locus of the multiplier spectrum morphism for polynomials of all degrees $d\geq 2$. Roughly speaking, non-injectivity implies intertwining, as it signifies the equivalence of polynomials or Ritt moves, except at isolated points and up to iteration. We investigate the relation between Ritt moves and the multiplier spectrum over arithmetic progressions.
Key words and phrases:
Multiplier spectrum, moduli space of polynomials, intertwined polynomials, generalized Lattès maps, periodic points.
2020 Mathematics Subject Classification:
Primary 37P35; Secondary 37P05, 37P45
1. Introduction
The aim of this paper is to study the multipliers of periodic points for a rational map $f$ on $\mathbb{P}^{1}(\mathbb{C})$ of degree at least $2$, concentrating on the case when $f$ is a polynomial.
1.1. The multiplier spectrum morphism on rational maps
We recall some general definitions, constructions, and results for rational maps (not just for polynomials).
Let $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$. For a $f$-periodic point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ with exact period $n$, we set $n_{f}(z_{0}):=n$ to denote this exact period. The multiplier $\rho_{f}(z_{0})$ of $f$ at $z_{0}$ is defined as the differential $df^{\circ n}(z_{0})\in\mathbb{C}$. (Throughout this thesis, the symbol $\circ$ is always used to denote composition of rational maps, while its absence indicates multiplication between rational maps.) We write $n(z_{0})=n_{f}(z_{0})$ and $\rho(z_{0})=\rho_{f}(z_{0})$ for simplicity, when the map $f$ is clear. The length of $f$ at $z_{0}$ is defined as the modulus $\lvert\rho_{f}(z_{0})\rvert$. The characteristic exponent of $f$ at $z_{0}$ is defined as $\chi_{f}(z_{0}):=n^{-1}\log\lvert\rho_{f}(z_{0})\rvert$ (when $\rho_{f}(z_{0})\neq 0$). Clearly, the multiplier, length, and characteristic exponent are all invariant under conjugacy by Möbius transformations.
Let ${\rm Per}(f)(\mathbb{C})$ denote the set of all periodic points in $\mathbb{P}^{1}(\mathbb{C})$ of $f$, and define
${\rm Per}^{*}(f)(\mathbb{C}):=\{z_{0}\in{\rm Per}(f)(\mathbb{C}):\rho_{f}(z_{0})\neq 0\}.$
We write ${\rm Per}(f)$ and ${\rm Per}^{*}(f)$ for simplicity when the base field $\mathbb{C}$ is clear. We let ${\rm PrePer}(f)(\mathbb{C})$ (or ${\rm PrePer}(f)$) denote the set of all $f$-preperiodic points in $\mathbb{P}^{1}(\mathbb{C})$.
We identify ${\rm PGL}_{2}(\mathbb{C})$ with the group of rational maps $g\in\mathbb{C}(z)$ of degree $1$ (i.e., the automorphism group of $\mathbb{P}^{1}$ over $\mathbb{C}$).
For an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $h$ is semi-conjugate to $g$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$, and we write $g\geq h$ (or $g\geq_{\pi}h$ when $\pi$ is specified).
In complex dynamics, the exceptional maps defined below are often regarded as special examples among all rational maps of degree $\geq 2$. These are rational maps related to algebraic groups that exhibit special dynamical properties.
Definition 1.1.
Let $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be an endomorphism over $\mathbb{C}$ of degree $d\geq 2$.
  • The map $f$ is called Lattès if there exists an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi\geq f$. Furthermore, the map $f$ is called flexible Lattès if there exist an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that $n]\geq_{\pi}f$, where $[n]$ denotes the multiplication-by-$n$ map on $E$ and $\pi:E\to\mathbb{P}^{1}$ is the quotient map modulo $\{\pm 1\}$. (Note that in this case we must have $d=n^{2}$, a perfect square.) A non-flexible Lattès map is called rigid Lattès.
  • We say that $f$ is of monomial type if it is semi-conjugate to the power map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$. It is known that $f$ is of monomial type if and only if it is conjugate to $z^{\pm d}$ or $\pm T_{d}(z)$, where $T_{d}(z)$ is the Chebyshev polynomial of degree $d$; see [Mil06].
  • $f$ is called exceptional if it is Lattès or of monomial type.
These definitions depend only on the conjugacy class of $f$. It is well-known that $f$ is exceptional if and only if some (hence every) iterate $f^{\circ k}$ is exceptional ($k\in\mathbb{Z}_{>0}$).
Fix an integer $d\geq 2$. Let ${\rm Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$, which is a smooth affine variety of dimension $2d+1$ [Sil12]. Let ${\rm FL}_{d}\subseteq{\rm Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}$. The group ${\rm PGL}_{2}={\rm Aut}(\mathbb{P}^{1})$ acts on ${\rm Rat}_{d}$ by conjugation. The geometric quotient
$\mathcal{M}_{d}:={\rm Rat}_{d}/{\rm PGL}_{2}={\rm Spec}(\mathcal{O}({\rm Rat}_{d})^{{\rm PGL}_{2}})$
is the (coarse) moduli space of endomorphisms of degree $d$ [Sil12], which is an affine variety of dimension 2d22d-2 [Sil07, Theorem 4.36(c)]. Let $\Psi:{\rm Rat}_{d}\to\mathcal{M}_{d}$ be the quotient morphism. For $f\in{\rm Rat}_{d}(\mathbb{C})$, we also denote its conjugacy class by $\Psi(f)=[f]$. Set ${\rm FL}_{d}]:=\Psi({\rm FL}_{d})$. Note that ${\rm FL}_{d}$ is non-empty if and only if $d$ is a perfect square, and in this case ${\rm FL}_{d}]$ is equidimensional of dimension $1$. The above construction works over any algebraically closed field of characteristic 0 and commutes with base changes.
Fix $n\in\mathbb{Z}_{>0}$. Let ${\rm Per}_{n}(f)={\rm Per}_{n}(f)(\mathbb{C})$ be the multi-set [^1] A multi-set is a set except allowing multiple instances for each of its elements. The number of the instances of an element is called its multiplicity. For example, $\{a,a,b,c,c,c\}$ is a multi-set of cardinality $6$, with multiplicities $2,1,3$ for $a,b,c$, respectively. of fixed points of $f^{\circ n}$. This multi-set has cardinality $d^{n}+1=:N_{d,n}$. Using elementary symmetric polynomials, the multipliers of the fixed points of $f^{\circ n}$ define a point
$S_{n}(f)=(\sigma_{1,n}(f),\dots,\sigma_{N_{d,n},n}(f))\in\mathbb{C}^{N_{d,n}},$
where $\sigma_{j,n}(f)$ is the $j$-th elementary symmetric polynomial in the $N_{d,n}$-tuple $\{\rho_{f^{\circ n}}(z)\}_{z\in{\rm Per}_{n}(f)}$ for $1\leq j\leq N_{d,n}$. We define the multiplier spectrum of $f$ to be the sequence $S(f)=(S_{n}(f))_{n=1}^{\infty}$. It is clear that $S_{n}$ takes the same value within a conjugacy class of rational maps and thus defines a morphism $S_{n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{A}^{N_{d,n}}(\mathbb{C})$, called the multiplier spectrum morphism of level $n$. We define the multiplier spectrum morphism up to level $n$ as
$\tau_{d,n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,n}}(\mathbb{C}),[f]\mapsto(S_{1}([f]),\dots,S_{n}([f])).$
For $f\in{\rm Rat}_{d}(\mathbb{C})$, we also write $\tau_{d,n}(f):=\tau_{d,n}([f])$.
For $n\in\mathbb{Z}_{>0}$, set $R_{d,n}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})^{2}\mid\tau_{d,n}([f])=\tau_{d,n}([g])\}$. Then $(R_{d,n})_{n=1}^{\infty}$ forms a decreasing sequence of Zariski closed subsets of $\mathcal{M}_{d}(\mathbb{C})^{2}$, which stabilizes by noetherianity. Let $m_{d}\geq 1$ be minimal such that $\cap_{n=1}^{\infty}R_{d,n}=R_{d,m_{d}}$. Consequently, for all $g\in{\rm Rat}_{d}(\mathbb{C}), ff$ and $g$ have the same multiplier spectrum if and only if $\tau_{d,m_{d}}(f)=\tau_{d,m_{d}}(g)$, i.e., they have the same multiplier spectrum up to level $m_{d}$. We define $\tau_{d}:=\tau_{d,m_{d}}$, called the multiplier spectrum morphism on $\mathcal{M}_{d}$ (or ${\rm Rat}_{d}$). It is well-known that rational maps within an irreducible component of ${\rm FL}_{d}(\mathbb{C})$ share the same multiplier spectrum.
The following remarkable theorem of McMullen [McM87] establishes that, with the exception of flexible Lattès maps, the multiplier spectrum determines the conjugacy class of rational maps up to only finitely many possibilities.
Theorem 1.2 (McMullen).
For every integer $d\geq 2$, the morphism
$\tau_{d}:\mathcal{M}_{d}(\mathbb{C})\setminus\Psi({\rm FL}_{d}(\mathbb{C}))\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
is quasi-finite.
Recently, Ji and Xie proved a significant generalization [JX25, Theorem 1.3] of McMullen’s theorem:
Theorem 1.3 (Ji and Xie).
For every integer $d\geq 2$, the morphism
$\tau_{d}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
is generically injective; that is, there exist a non-empty Zariski open subset $U$ of $\mathcal{M}_{d}(\mathbb{C})$ and a Zariski open subset $W$ of the Zariski closure of $\tau_{d}(U)$ with $\tau_{d}^{-1}(W)=U$ such that $\tau_{d}:U\to W$ is a finite morphism of degree $1$.
Similarly, by replacing the multiplier with its modulus and $\mathbb{C}$ with $\mathbb{R}$ in the above definitions, we obtain for each $n\in\mathbb{Z}_{>0}$ the length spectrum map $L_{n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}^{N_{d,n}}$ of level $n$. The length spectrum of $f$ is defined as the sequence $L(f)=(L_{n}(f))_{n=1}^{\infty}$. We define the length spectrum map up to level $n$ as
$\eta_{d,n}:\mathcal{M}_{d}(\mathbb{C})\to\mathbb{R}^{N_{d,1}}\times\cdots\times\mathbb{R}^{N_{d,n}},[f]\mapsto(L_{1}([f]),\dots,L_{n}([f])).$
(For the time being, we treat these maps merely as maps between sets.)
1.2. Statement of the main results
Fix an integer $d\geq 2$. By restriction, we obtain the multiplier spectrum morphism for polynomials. Let ${\rm Poly}^{d}\subset{\rm Rat}_{d}$ be the closed subvariety of polynomials of degree $d$, and let ${\rm MPoly}^{d}:={\rm Poly}^{d}/{\rm Aff}$ be the moduli space of polynomials of degree $d$, see [FG22]. Here $\rm Aff$} is the group of automorphisms of $\mathbb{A}^{1}$ consisting of all linear polynomials, which acts on ${\rm Poly}^{d}$ by conjugation. It is clear that for all $g\in{\rm Poly}^{d}(\mathbb{C}), ff$ and $g$ are conjugate by a Möbius transformation $\tau\in{\rm PGL}_{2}(\mathbb{C})$ if and only if they are conjugate by a linear polynomial $\sigma\in{\rm Aff}(\mathbb{C})$. Let $\Psi:{\rm Poly}^{d}\to{\rm MPoly}^{d}$ still denote the quotient morphism. For every $f\in{\rm Poly}^{d}(\mathbb{C})$, we also let $\Psi(f)=[f]\in{\rm MPoly}^{d}(\mathbb{C})$ denote its conjugacy class. We can view ${\rm MPoly}^{d}\subset\mathcal{M}_{d}$, which is compatible with $\Psi$ and ${\rm Poly}^{d}\subset{\rm Rat}_{d}$. For every integer $n\geq 1$, denote the restriction of $\tau_{d,n}$ (resp. $\tau_{d}$) to the moduli space of polynomials ${\rm MPoly}^{d}(\mathbb{C})$ by $\tilde{\tau}_{d,n}$ (resp. $\tilde{\tau}_{d}$). Similarly, we obtain $\tilde{\eta}_{d,n}$ for the length spectrum.
Generic injectivity
Ji and Xie also prove the following polynomial version of Theorem 1.3, which is also proved by Huguin [Hug24] independently using completely different methods.
Theorem 1.4 ([JX25, Theorem 1.4] and [Hug24, Theorem C]).
For every integer $d\geq 2$, the morphism
$\tilde{\tau}_{d}:{\rm MPoly}^{d}(\mathbb{C})\to\mathbb{A}^{N_{d,1}}(\mathbb{C})\times\cdots\times\mathbb{A}^{N_{d,m_{d}}}(\mathbb{C})$
is generically injective.
In [JX25], the proof of Theorem 1.4 is similar to the proof of Theorem 1.3 in the same paper. Specifically, the proof of Theorem 1.3 relies on two key ingredients: a variant of the Dynamical André-Oort (DAO) conjecture for curves [JX25, Theorem 3.4] and a result [JX25, Theorem 3.3] based on Pakovich’s work [Pak25a]. In contrast, the proof of Theorem 1.4 replaces the latter ingredient with [FG22, Theorems 3.51 and 3.52]. Huguin’s proof [Hug24] actually shows that $\tilde{\tau}_{d,2}$ is generically injective for every $d\geq 2$, and it is based on intricate local computations.
The result [JX25, Theorem 3.3] involves a specific class of rational maps, namely simple rational maps. Recall that a rational map $G\in\mathbb{C}(z)$ of degree $m\geq 2$ is called simple if $G$ has exactly 2m22m-2 distinct critical values in $\mathbb{P}^{1}(\mathbb{C})$. However, a polynomial of degree $d\geq 3$ is never simple, due to the totally invariant fixed point $\infty$. To overcome this problem, for a polynomial $F\in\mathbb{C}[z]$ of degree $m\geq 2$, we call $F$ pre-simple if it has exactly m1m-1 distinct critical values in $\mathbb{C}$. This class of pre-simple polynomials serves as an analogue of simple rational maps in the polynomial case. We find that Pakovich’s methods [Pak25a] are also applicable to pre-simple polynomials. In §3.1, we present the following polynomial version of [JX25, Theorem 3.3] (for the case $d\geq 4$), which is essentially due to Pakovich [Pak25a]:
Theorem 1.5.
For every integer $d\geq 4$, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that, for every $f\in U$, the following hold:
  1. (1)
    $f$ is pre-simple;
  2. (2)
    for every pre-simple $g\in{\rm Poly}^{d}(\mathbb{C})$, if $f$ and $g$ are intertwined, then $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Recall that two polynomials $F,G\in\mathbb{C}[z]$ of degree $\geq 2$ are called intertwined if there exists a (possibly reducible) algebraic curve $Z\subset(\mathbb{A}^{1})^{2}$ whose projections to both axes are onto, and $Z$ is invariant under the endomorphism $(F,G):(\mathbb{A}^{1})^{2}\to(\mathbb{A}^{1})^{2}$.
Recently, Pakovich [Pak25b] has studied general rational maps of degree $2\leq d\leq 3$ in detail. Consequently, one can prove Theorem 1.3 for this case using the results in [Pak25b] as part of the argument. The philosophy is that a general rational map of degree $d$ has “good” properties regarding the decompositions of its iterates. However, the simplicity condition is not sufficient when $2\leq d\leq 3$, unlike the situation in [Pak25a] where $d\geq 4$.
Following Pakovich [Pak25b], we study general polynomial maps of degree $2\leq d\leq 3$ in §3.2 and show the following polynomial version of [JX25, Theorem 3.3] for the case where $2\leq d\leq 3$:
Proposition 1.6.
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$ with $f$ and $g$ intertwined, we have $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Proposition 1.6 is also a direct corollary of [FG22, Theorems 3.51 and 3.52] (see also [GNY19, Theorem 1.4]). We include this to present different approaches.
In §3.3, we provide a detailed proof of Theorem 1.4 using Theorem 1.5 and Proposition 1.6, which is essentially due to Pakovich [Pak25a, Pak25b] and Ji-Xie [JX25].
The intermediate results presented in §3 may be useful for studying the dynamics of one-variable polynomials and their products. For instance, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $(\mathbb{P}^{1})^{2}$ with all factors of degree $d\geq 2$ (see Theorem 3.33).
The non-injective locus
After establishing the generic injectivity of the multiplier spectrum morphism, a natural further direction is to study the non-injective locus of $\tau_{d}$ (resp. $\tilde{\tau}_{d}$), i.e., the locus
$\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})^{2}\mid\tau_{d,m_{d}}([f])=\tau_{d,m_{d}}([g])\}\setminus\Delta,$
or its intersection with ${\rm MPoly}^{d}(\mathbb{C})^{2}$, where $\Delta$ is the diagonal in $\mathcal{M}_{d}(\mathbb{C})^{2}$.
There are two known cases of non-conjugate rational maps with the same multiplier spectrum. For a non-constant rational map $f$, if $f=h_{1}\circ h_{2}$ in $\mathbb{C}(z)$, then $h:=h_{2}\circ h_{1}$ is called an elementary transformation of $f$. Two non-constant rational maps $f$ and $g$ are called equivalent, written $f\sim g$, if there exists a finite chain of elementary transformations between $f$ and $g$. Clearly, equivalence defines an equivalence relation for non-constant rational maps, which is coarser than conjugacy (i.e., conjugacy implies equivalence), so we can talk about the equivalence of two conjugacy classes. An easy lemma of Pakovich [Pak19, Lemma 2.1] shows that if $f$ and $g$ are equivalent (of degree $\geq 2$), then they have the same multiplier spectrum. The other known case of non-injectivity comes from exceptional maps. Silverman [Sil07, Theorem 6.62] showed that $\tau_{d}$ is not injective at $[f]$ for some rigid Lattès maps $f$. Pakovich asked [Pak19, Problem 3.1] whether these are the only obstructions to the injectivity of $\tau_{d}$. Ji and Xie [JX25, Conjecture 1.5] conjectured that these are the only obstructions to the injectivity of $\tau_{d}$:
Conjecture 1.7.
Let $f$ and $g$ be non-conjugate rational maps of degree $d\geq 2$ with the same multiplier spectrum. Then one of the following holds:
  1. (i)
    $f$ and $g$ are Lattès maps;
  2. (ii)
    $f\sim g$.
When $f$ and $g$ are polynomials, only case ii in Conjecture 1.7 can occur. Note that if the degree $d\geq 2$ is a prime number, then two rational maps $f$ and $g$ of degree $2\leq d\leq 3$ are equivalent if and only if they are conjugate. Consequently, Conjecture 1.7 predicts that for every prime $p\in\mathbb{Q}$, the multiplier spectrum morphism $\tilde{\tau}_{p}$ is injective on ${\rm MPoly}^{p}(\mathbb{C})$. On the other hand, an observation of Huguin [Hug24, Corollary 85] shows that $\tilde{\tau}_{d}$ is not injective for every composite $d\geq 2$.
If the degree $d$ is small, then it is possible to analyze the non-injective locus of $\tilde{\tau}_{d}$ (indeed, $\tilde{\tau}_{d,m}$ with small $m$) using explicit computations. For the prime degree $2\leq d\leq 3$, Huguin [Hug24, Theorem C] showed that $\tilde{\tau}_{d,1}$ is injective on the moduli space ${\rm MPoly}^{d}(\mathbb{C})$. (The injectivity of $\tilde{\tau}_{2,1}$ (in fact, $\tau_{2,1}$) was first proved by Milnor [Mil93].) For the composite degree $d=4$, a result [Hug24, Proposition 86] of Huguin implies that for every $(f,g)\in{\rm Poly}^{4}(\mathbb{C})^{2}$, $\tilde{\tau}_{4}(f)=\tilde{\tau}_{4}(g)$ if and only if $\tilde{\tau}_{4,2}(f)=\tilde{\tau}_{4,2}(g)$ if and only if $[f]=[g]$ in ${\rm MPoly}^{4}(\mathbb{C})$ or $(f,g)=(h_{1}\circ h_{2},h_{2}\circ h_{1})$ for some $(h_{1},h_{2})\in{\rm Poly}^{2}(\mathbb{C})^{2}$ (in particular, $f\sim g$), whose proof relies on computations using a computer.
We intend to investigate the non-injective locus of the multiplier spectrum morphism for polynomials. In §4, for all $d\geq 2$, we provide a first description of the non-injective locus of $\tilde{\tau}_{d}$, i.e., the locus
${\rm PNI}_{d}:=\{([f],[g])\in{\rm MPoly}^{d}(\mathbb{C})^{2}\mid\tau_{d,m_{d}}([f])=\tau_{d,m_{d}}([g]),[f]\neq[g]\},$
which is a starting point of Conjecture 1.7 for the polynomial case:
Theorem 1.8.
Let $d\geq 2$ be an integer, and let $C\subset{\rm MPoly}^{d}\times{\rm MPoly}^{d}$ be an irreducible curve defined over $\overline{\mathbb{Q}}$ such that the projection of $C$ to each factor ${\rm MPoly}^{d}$ is non-constant. Assume $C(\mathbb{C})\setminus{\rm PNI}_{d}$ is finite. Then for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined.
In fact, we prove a more precise version:
Theorem 1.9.
Let $\phi_{1},\phi_{2}:C\times\mathbb{P}^{1}\to C\times\mathbb{P}^{1}$ be two non-isotrivial (algebraic) families over $\overline{\mathbb{Q}}$ of degree-$d$ ($d\geq 2$) polynomials, parameterized by an affine irreducible curve $C$ over $\overline{\mathbb{Q}}$. Denote the induced morphisms also by $\phi_{1},\phi_{2}:C\to{\rm MPoly}^{d}$. Assume $(\phi_{1}(t),\phi_{2}(t))\in{\rm PNI}_{d}$ for all but finitely many $t\in C(\mathbb{C})$. (Here non-isotriviality means that $\phi_{j}(C)$ is not a single point in ${\rm MPoly}^{d}$ for $j=1,2$.)
Then there exists a finite subset $S$ of $C(\mathbb{C})$ such that, for every $t\in C(\mathbb{C})\setminus S$, one of the following conditions holds:
  1. (1)
    there exists $N\in\mathbb{Z}_{>0}$ such that $\phi_{1}(t)^{\circ N}\sim\phi_{2}(t)^{\circ N}$;
  2. (2)
    there exist $N,k_{1},k_{2},l\in\mathbb{Z}_{>0}$ and $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$, with k_1k_2,gcd(d,k)=1,l1 (mod k),V(0)0 such that ϕ_1(t)^Nz^lV(z^k_1)^k_1^and ϕ_2(t)^Nz^lV(z^k_2)^k_2^, where $k:={\rm lcm}(k_{1},k_{2})$ denotes the least common multiple of $k_{1}$ and $k_{2}$, and $k_{j}^{\prime}=k/k_{j}$ for $j=1,2$.
Remark 1.10.
In Theorem 1.9 $(2)$, we have
$(\phi_{1}(t)^{\circ N})_{\rm min}\sim(\phi_{2}(t)^{\circ N})_{\rm min},$
see Remark 4.4.
In (2), if $k_{j}^{\prime}>1$ ($j=1,2$), then $z^{l}\cdot V(z^{k_{j}})^{k_{j}^{\prime}}$ is a generalized Lattès map in the sense of Pakovich [Pak20c] (see Definition 2.3 and Proposition 2.4). Note that at least one $k_{j}^{\prime}>1$ since $k_{1}\neq k_{2}$.
Remark 1.11.
In Theorem 1.9, we can choose NN in (1) or (2) to have an upper bound depending only on $\mathbb{Q}(\phi_{1},\phi_{2})$ and $d$, see Remark 4.5.
The proof of Theorem 1.9 is divided into two steps. The first step is as follows. After a base change, we may assume that all the critical points are marked for $\phi_{1}$ and $\phi_{2}$. It is well-known that post-critically finite (PCF) parameters do not form families in ${\rm MPoly}^{d}$. We argue that the multiplier spectrum determines the periodicity of critical points; see Proposition 4.3 for a precise statement. Then from $\phi_{1}$ and $\phi_{2}$ we can construct two entangled active dynamical pairs in the sense of Favre-Gauthier [FG22], still denoted by $\phi_{1}$ and $\phi_{2}$ with parameters in $C$. Applying a theorem of Favre-Gauthier (Theorem 4.1), we see that both $\phi_{1}^{\circ N}$ and $\phi_{2}^{\circ N}$ are semi-conjugate to a common family $R$ of polynomials over a non-empty Zariski open subset $U$ of $C$, for some integer $N\geq 1$. In the second step, we need some results of Pakovich on generalized Lattès maps [Pak20c]. Set $S=(C\setminus U)(\mathbb{C})$, which is a finite set. Fix $t\in U(\mathbb{C})$. Assume that (1) does not hold for $t$. Then [Pak20c, Theorem 2.10] implies that $R(t)$ is a generalized Lattès map. By the classification of generalized Lattès maps that are polynomials (Proposition 2.4) and some elementary arguments, we conclude that (2) holds for $t$, after possibly further iteration.
1.3. Organization of the article
In §2, we provide a brief introduction to orbifolds and generalized Lattès maps based on [Pak20c], for the needs of §3 and §4.
In §3, we explore the generic injectivity of the multiplier spectrum morphism for polynomials. In §3.1, we introduce the notion of pre-simple polynomials and study their properties, essentially done by Pakovich [Pak25a]. Then we deduce Theorem 1.5. In §3.2, we study general polynomials of degree $d\in\{2,3\}$ and obtain Proposition 1.6, following the ideas in [Pak25b]. In §3.3, we deduce Theorem 1.4 from Theorem 1.5 and Proposition 1.6, following Ji-Xie [JX25]. Finally, in §3.4, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $(\mathbb{P}^{1})^{2}$ with all factors of degree $d\geq 2$ (Theorem 3.33).
In §4.1, we recall the notion of critically marked polynomial and state a theorem of Favre-Gauthier (Theorem 4.1), which is essential for the proof of Theorem 1.9. Then we proceed to complete the proofs of Theorem 1.9 and Theorem 1.8 in §4.2. Subsection 4.3 is devoted to the study of the situation related to Theorem 1.9 (2). We consider pairs of polynomials of the form $(z^{r}\cdot R(z)^{k},z^{r}\cdot R(z^{k}))$, called Ritt moves in this article. We show that under specific conditions $z^{r}\cdot R(z)^{k}$ and $z^{r}\cdot R(z^{k})$ have the same multiplier spectrum over an arithmetic progression (Theorem 4.6) and that they always have the same set of multipliers up to powers (Proposition 4.11).
In §5, we present many problems and questions for future study. In §5.1, multiplier spectrum over arithmetic progressions and the related notion of stable multiplier spectrum are considered. In §5.2, we present several directions to generalize Theorem 1.9.
Acknowledgement
The author would like to express his great gratitude to his advisor, Professor Junyi Xie, for the constant encouragement. The author is supported by NSFC Grant (No. 12271007). The author would like to thank Valentin Huguin for helpful comments on the first version of this paper.
2. Orbifolds and generalized Lattès maps
This section collects definitions and results concerning orbifolds and generalized Lattès maps, as needed for our purposes. The theory of generalized Lattès maps (based on orbifolds) was primarily developed by Pakovich; see [Pak20c, Pak16, Pak18, Pak20a, Pak20b, Pak23]. Here, we restrict our attention to orbifolds on the compact Riemann surface $\mathbb{P}^{1}(\mathbb{C})$, which suffices for our needs.
Definition 2.1.
An orbifold $\mathcal{O}$ on $\mathbb{P}^{1}(\mathbb{C})=\widehat{\mathbb{C}}$ is defined by a ramification function $\nu:\widehat{\mathbb{C}}\to\mathbb{Z}_{>0}$ such that $\nu^{-1}(\left[2,\infty\right[)$ is discrete. We only consider good orbifolds, i.e., those satisfying $\nu^{-1}(\left[2,\infty\right[)\neq 1$ and, if $\nu^{-1}(\left[2,\infty\right[)=2$, then $\nu(\widehat{\mathbb{C}})=2$. For an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with ramification function $\nu$, its Euler characteristic is defined as
$\chi(\mathcal{O}):=2+\sum_{z\in\widehat{\mathbb{C}}}\left(\frac{1}{\nu(z)}-1\right),$
the set of singular points is $c(\mathcal{O})=\{z\in\widehat{\mathbb{C}}:\nu(z)>1\}$, and the signature is the multi-set $\nu(\mathcal{O})=\{\nu(z):z\in c(\mathcal{O})\}$ of cardinality $c(\mathcal{O})$.
Definition 2.2.
Let $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ be two orbifolds on $\widehat{\mathbb{C}}$ with ramification functions $\nu_{1}$ and $\nu_{2}$, respectively, and let $f:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be a non-constant rational map.
  1. (1)
    We say $f:\mathcal{O}_{1}\to\mathcal{O}_{2}$ is a covering map between orbifolds if for every $z\in\widehat{\mathbb{C}}$, we have ν_2(f(z))=ν_1(z)⋅deg_z(f), where $\deg_{z}(f)$ denotes the local degree of $f$ at $z$.
  2. (2)
    We say $f:\mathcal{O}_{1}\to\mathcal{O}_{2}$ is a minimal holomorphic map between orbifolds if for every $z\in\widehat{\mathbb{C}}$, we have ν_2(f(z))=ν_1(z)⋅gcd(deg_z(f),ν_2(f(z))).
Clearly, a covering map between orbifolds is also a minimal holomorphic map.
An equivalent description of Lattès maps is well-known: a rational map $A:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $\geq 2$ is a Lattès map if and only if there exists an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with a non-constant ramification function such that $A:\mathcal{O}\to\mathcal{O}$ is a covering map between orbifolds. This motivates the following definition:
Definition 2.3.
A rational map $A:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $\geq 2$ is called a generalized Lattès map if there exists an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with a non-constant ramification function such that $A:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map between orbifolds.
From the definition, it is clear that all exceptional rational maps of degree $\geq 2$ are generalized Lattès maps. From [Pak20c, Proof of Theorem 7.2], we obtain the following complete description of polynomials that are generalized Lattès maps:
Proposition 2.4.
Let $f\in\mathbb{C}[z]$ be a polynomial of degree at least $2$.
  1. (1)
    If $f$ is exceptional, then $f$ is a generalized Lattès map.
  2. (2)
    Suppose $f$ is non-exceptional. If $\mathcal{O}$ is an orbifold on $\widehat{\mathbb{C}}$ with a non-constant ramification function such that $f:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map between orbifolds, then $\nu(\mathcal{O})=\{n,n\}$ for some integer $n\geq 2$, and there exist a positive integer $r$ with $\gcd(r,n)=1$ and a non-constant polynomial $R\in\mathbb{C}[z]\setminus\mathbb{C}$ such that $f$ is conjugate to $z^{r}\cdot R(z)^{n}$. In fact, $f$ is a generalized Lattès map if and only if there exist integers $r\geq 1$ and $n\geq 2$ with $\gcd(r,n)=1$, and a polynomial $R\in\mathbb{C}[z]$ such that $f$ is conjugate to $z^{r}\cdot R(z)^{n}$.
Using [Pak20c, Theorem 1.1], we easily obtain the following observation:
Lemma 2.5.
Let $f,g\in\mathbb{C}(z)\setminus\mathbb{C}$ be rational maps of degree $\geq 2$ such that $f\geq g$. If $f$ is a generalized Lattès map, then so is $g$.
Two natural orbifolds are associated with a rational map:
Definition 2.6.
Let $f:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be a non-constant rational map.
  1. (i)
    The orbifold $\mathcal{O}_{2}^{f}$ is defined by the ramification function ν_2(z)=lcm{deg_w(f):w∈f^-1(z)}, z∈^C.
  2. (ii)
    The orbifold $\mathcal{O}_{1}^{f}$ is defined by the ramification function ν_1(z)=ν_2(f(z))/deg_z(f), z∈^C.
It is straightforward to verify that a non-constant rational map $f:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ is a covering map $f:\mathcal{O}_{1}^{f}\to\mathcal{O}_{2}^{f}$ between orbifolds.
Let $f:\mathbb{P}^{1}_{\mathbb{C}}\to\mathbb{P}^{1}_{\mathbb{C}}$ be a non-constant rational map over $\mathbb{C}$, i.e., $f\in\mathbb{C}(z)\setminus\mathbb{C}$. A rational map $f_{0}\in\mathbb{C}(z)$ is called a compositional right factor of $f$ if there exists $f_{1}\in\mathbb{C}(z)$ such that $f=f_{1}\circ f_{0}$. Similarly, we define compositional left factors. Pakovich [Pak20c] introduced the notion of good solutions for certain equations, which includes the following special case:
Definition 2.7.
Consider the functional equation
(2.1) $f\circ p=g\circ q,$
where $f,p,g,q:\mathbb{P}^{1}_{\mathbb{C}}\to\mathbb{P}^{1}_{\mathbb{C}}$ are non-constant rational maps over $\mathbb{C}$. A solution $(f,p,g,q)$ of (2.1) is called good if the fiber product of $f$ and $g$ consists of a unique component, and $p$ and $q$ share no common compositional right factor of degree at least $2$.
Pakovich [Pak20c, Theorem 5.1] described all good solutions of (2.1) with $p(z)=g(z)=z^{n}$ as follows:
Theorem 2.8.
Let $n\geq 2$ be an integer and $A,F\in\mathbb{C}(z)$ be rational maps of degree at least $2$ such that $A\circ z^{n}=z^{n}\circ F$. Then the following are equivalent:
  1. (1)
    The solution $(A,z^{n},z^{n},F)$ of (2.1) is good.
  2. (2)
    There exist $R\in\mathbb{C}(z)$ and $r\in\mathbb{Z}_{>0}$ with $\gcd(r,n)=1$ such that A(z)=z^rR(z)^n and F(z)=z^rR(z^n).
Remark 2.9.
When both $A$ and $F$ are polynomials, we may take $R$ in (2) to be a polynomial as well.
Pakovich [Pak20c] showed that rational maps that are not generalized Lattès maps exhibit “better” behavior than generalized Lattès maps. The following theorem, a special case of [Pak20c, Theorem 3.2], illustrates this.
Theorem 2.10 (Pakovich).
Let $A,B,\pi:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be rational maps of degree at least $2$ with $A\circ\pi=\pi\circ B$. Then one of the following holds:
  1. (1)
    $B\sim A$;
  2. (2)
    There exist non-constant rational maps $\psi,\pi_{0},B_{0}$ satisfying:
    • AA is a generalized Lattès map;
    • $\pi=\pi_{0}\circ\psi$ and $\deg(\pi_{0})\geq 2$;
    • $B_{0}\circ\psi=\psi\circ B$, $A\circ\pi_{0}=\pi_{0}\circ B_{0}$, and $B\sim B_{0}$;
    • $(A,\pi_{0},\pi_{0},B_{0})$ is a good solution of (2.1);
    • $A:\mathcal{O}_{2}^{\pi_{0}}\to\mathcal{O}_{2}^{\pi_{0}}$ and $B_{0}:\mathcal{O}_{1}^{\pi_{0}}\to\mathcal{O}_{1}^{\pi_{0}}$ are minimal holomorphic maps between orbifolds;
    • There exists $s\in\mathbb{Z}_{>0}$ such that $\psi$ is a compositional right factor of $B^{\circ s}$ and a compositional left factor of $B_{0}^{\circ s}$.
3. The multiplier spectrum morphism is generically injective for polynomials
3.1. Pre-simple polynomials and the case $d\geq 4$
A particular class of rational maps, namely simple rational maps, plays a role in the proof of Theorem 1.3 in [JX25] (see [JX25, Theorem 3.3]). This notion was introduced by Pakovich [Pak25a]. A rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is simple if it has exactly 2d22d-2 distinct critical values in $\mathbb{P}^{1}(\mathbb{C})$. By the Riemann–Hurwitz formula, every $f\in\mathbb{C}(z)$ of degree $d$ has exactly 2d22d-2 critical points counting with multiplicity. Hence, a simple rational map is a rational map with all critical points of multiplicity $1$ such that no two distinct critical points map to the same value.
For rational maps, simplicity is a natural notion since simple rational maps of degree $d\geq 2$ form a non-empty Zariski open subset of ${\rm Rat}_{d}(\mathbb{C})$. However, every polynomial $f\in\mathbb{C}[z]$ of degree $d\geq 2$ has a critical point of multiplicity d1d-1 at $\infty$. Thus, $f$ is simple if and only if $d=2$. The notion of simplicity is not useful for polynomials. To overcome this, we introduce a modified notion for polynomials:
Definition 3.1.
A polynomial $f(z)\in\mathbb{C}[z]$ of degree $d\geq 2$ is called pre-simple if it has exactly d1d-1 distinct critical values in $\mathbb{C}$.
We show that most statements in [Pak25a] concerning simplicity for rational maps also hold for pre-simplicity for polynomials. In particular, we obtain a polynomial version (Theorem 3.23) of [JX25, Theorem 3.3], which yields a proof of Theorem 1.4 for $d\geq 4$. Most proofs in this section are minor modifications of those in [Pak25a]; we provide complete details for the reader’s convenience.
We first recall some notions about decomposing rational maps. A decomposition of a rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is a representation $f=f_{r}\circ\cdots\circ f_{1}$, where each $f_{i}$ has degree $\geq 2$. Two decompositions $f=f_{r}\circ\cdots\circ f_{1}$ and $f=g_{l}\circ\cdots\circ g_{1}$ are equivalent if either $l=r=1$ and $f_{1}=g_{1}$, or $l=r\geq 2$ and there exist Möbius transformations $\nu_{1},\dots,\nu_{r-1}\in{\rm PGL}_{2}(\mathbb{C})$ such that $f_{r}=g_{r}\circ\nu_{r-1}$, $f_{j}=\nu_{j}^{-1}\circ g_{j}\circ\nu_{j-1}$ for $1<j<r$, and $f_{1}=\nu_{1}^{-1}\circ g_{1}$. A rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is indecomposable if $f=f_{2}\circ f_{1}$ implies $\deg(f_{1})=1$ or $\deg(f_{2})=1$.
Proposition 3.2.
A pre-simple polynomial $f$ of degree $d\geq 2$ is indecomposable (as a rational map) and satisfies ${\rm Mon}(f)\cong S_{d}$, where $S_{d}$ is the permutation group on $\{1,2,\dots,d\}$ and ${\rm Mon}(f)$ is the monodromy group of $f$.
Proof.
Let $f\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $d\geq 2$. Suppose $f$ is decomposable, so $f=f_{1}\circ f_{2}$ with $f_{1},f_{2}\in\mathbb{C}(z)$ of degrees $m_{1},m_{2}\geq 2$, respectively. Then $d=m_{1}m_{2}$. Since $\infty$ is a totally ramified fixed point of $f=f_{1}\circ f_{2}$, the point $a=f_{2}(\infty)$ satisfies $f_{1}^{-1}(\infty)=\{a\}$ and $f_{2}^{-1}(a)=\{\infty\}$. After composing with suitable Möbius transformations, we may assume $a=\infty$ and both $f_{1}$ and $f_{2}$ are polynomials. The pre-simplicity of $f$ implies that the number of critical values of $f$ in $\mathbb{C}$ is N(f)=(2d2)(d1)=d1N(f)=(2d-2)-(d-1)=d-1. On the other hand, from $f^{\prime}=(f_{1}^{\prime}\circ f_{2})\cdot f_{2}^{\prime}$, we have
$m_{1}m_{2}-1=d-1=N(f)\leq N(f_{1})+N(f_{2})\leq m_{1}-1+m_{2}-1.$
Equivalently, $(m_{1}-1)(m_{2}-1)\leq 0$, contradicting $m_{1},m_{2}\geq 2$. Hence, $f$ is indecomposable.
Since $f$ is indecomposable, its monodromy group ${\rm Mon}(f)$ is primitive. Fix a critical value $c$ of $f$ in $\mathbb{C}$; the corresponding permutation in ${\rm Mon}(f)$ is a transposition. By [Wie64, Theorem 13.3], ${\rm Mon}(f)$ must be the full symmetric group, so ${\rm Mon}(f)\cong S_{d}$. ∎
We now introduce some algebraic curves associated with polynomials, following [Pak25a, §2].
Definition 3.3.
Let $F(z),H(z)\in\mathbb{C}[z]$ be polynomials of degrees $m,n\geq 1$, respectively. Define the algebraic curves $h_{F,H}$ and $h_{F}$ by
$h_{F,H}:F(x)-H(y)=0\quad\text{and}\quad h_{F}:\frac{F(x)-F(y)}{x-y}=0.$
The genera of these curves can be computed explicitly as follows; see [Fri74] or [Pak11].
Fact 3.4.
Let $F(z),H(z)\in\mathbb{C}[z]$ be polynomials of degrees $m,n\geq 1$, respectively, and let $S=\{z_{1},\dots,z_{r}\}$ be the set of all critical values of $F$ or $H$ in $\mathbb{C}$. For $1\leq j\leq r$, let $(a_{j,1},\dots,a_{j,p_{j}})$ and $(b_{j,1},\dots,b_{j,q_{j}})$ denote the multiplicities of $F$ and $H$ at the points in $F^{-1}(z_{j})$ and $H^{-1}(z_{j})$, respectively. Then the genera of $h_{F,H}$ and $h_{F}$ satisfy
$2-2g(h_{H,F})=\gcd(m,n)-(r-1)mn+\sum_{j=1}^{r}\sum_{s_{1}=1}^{p_{j}}\sum_{s_{2}=1}^{q_{j}}\gcd(a_{j,s_{1}},b_{j,s_{2}})$
and
$4-2g(h_{F})=m-(r-1)m^{2}+\sum_{j=1}^{r}\sum_{s_{1}=1}^{p_{j}}\sum_{s_{2}=1}^{p_{j}}\gcd(a_{j,s_{1}},a_{j,s_{2}}).$
The following two theorems are analogues of [Pak25a, Theorems 2.3 and 2.4].
Theorem 3.5.
Let $m\geq 4$ and $n\geq 2$ be integers, with $n\neq 2,4$ when $m=4$. Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m$ and $H\in\mathbb{C}[z]$ a polynomial of degree $n$. Assume the curve $h_{F,H}$ is irreducible. Then $g(h_{F,H})>0$. In particular, there exist no non-constant rational functions $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=H\circ Y$ in $\mathbb{C}(z)$.
Proof.
We use the notation from Fact 3.4. Let ${\rm critV}(F)$ and ${\rm critV}(H)$ denote the sets of critical values of $F$ and $H$ in $\mathbb{C}$, respectively.
For $i\in\{1,\dots,r\}$ with $z_{i}\notin{\rm critV}(F)$, we have $p_{i}=m$ and $(a_{i,1},\dots,a_{i,m})=(1,\dots,1)$, so
$\sum_{j_{1}=1}^{p_{i}}\sum_{j_{2}=1}^{q_{i}}\gcd(a_{i,j_{1}},b_{i,j_{2}})=\sum_{j_{1}=1}^{p_{i}}q_{i}=p_{i}q_{i}=mq_{i}.$
For $i\in\{1,\dots,r\}$ with $z_{i}\in{\rm critV}(F)$, since $F$ is pre-simple, we have $\deg_{z}(F)\in\{1,2\}$ for every $z\in\mathbb{C}$. Then $p_{i}=m-1$ and we may assume $(a_{i,1},\cdots,a_{i,m-2},a_{i,m-1})=(1,\dots,1,2)$. Let $l_{i}$ be the number of indices $j_{2}\in\{1,\dots q_{i}\}$ such that $b_{i,j_{2}}$ is even. Then
$\sum_{j_{1}=1}^{p_{i}}\sum_{j_{2}=1}^{q_{i}}\gcd(a_{i,j_{1}},b_{i,j_{2}})=(m-2)q_{i}+q_{i}+l_{i}=mq_{i}+(l_{i}-q_{i}).$
By the Riemann–Hurwitz formula,
$(2n-2)-(n-1)=\sum_{z\in\mathbb{C}}(\deg_{z}(H)-1)=\sum_{i=1}^{r}\sum_{j_{2}=1}^{q_{i}}(b_{i,j_{2}}-1)=rn-\sum_{i=1}^{r}q_{i},$
so $\sum_{i=1}^{r}q_{i}=(r-1)n+1$. Thus,
$\sum_{i=1}^{r}\sum_{j_{1}=1}^{p_{i}}\sum_{j_{2}=1}^{q_{i}}\gcd(a_{i,j_{1}},b_{i,j_{2}})=\sum_{i=1}^{r}mq_{i}+\Sigma=(r-1)mn+m+\Sigma,$
where $\Sigma:=\sum_{1\leq i\leq r,\,z_{i}\in{\rm critV}(F)}(l_{i}-q_{i})$. By Fact 3.4, the genus is
$g(h_{H,F})=\frac{1}{2}\cdot(2-m-\gcd(m,n)-\Sigma).$
Hence, $g(h_{H,F})=0$ if and only if $\Sigma=m+\gcd(m,n)-2$. Note that $\Sigma$ equals the number of $z\in\mathbb{C}$ such that $H(z)\in{\rm critV}(F)$ and $\deg_{z}(H)$ is odd.
For any finite subset $A\subset\mathbb{C}$, we have $n-1\geq\sum_{z\in H^{-1}(A)}(\deg_{z}(H)-1)$; hence $H^{-1}(A)\geq n(\#A)-(n-1)=n(-1+\#A)+1$, with equality if and only if ${\rm critV}(H)\subseteq S$. In particular, $H^{-1}({\rm critV}(F))\geq n(m-2)+1$, with equality if and only if ${\rm critV}(H)\subseteq{\rm critV}(F)$.
Suppose $g(h_{F,H})=0$. Then $\Sigma=m+\gcd(m,n)-2$, which implies
$\#H^{-1}({\rm critV}(F))$
$\leq$ $m+\gcd(m,n)-2+\left\lfloor\frac{n(m-1)-m-\gcd(m,n)+2}{2}\right\rfloor$
$=$ $\frac{1}{2}((m-1)(n+1)+\gcd(m,n)-1).$
Thus, $(m-1)(n+1)+\gcd(m,n)-1\geq 2n(m-2)+2$, or equivalently, $mn+4\leq 3n+m+\gcd(m,n)$. Since $\gcd(m,n)\leq n$, we get $mn+4\leq 4n+m$, so $(m-4)(n-1)\leq 0$. Given $n\geq 2$ and $m\geq 4$, we must have $m=4$. Then $n\neq 2,4$ by assumption. Now $mn+4\leq 3n+m+\gcd(m,n)$ becomes $n\leq\gcd(4,n)$, contradicting $n\neq 2,4$. Therefore, $g(h_{H,G})>0$.
Now suppose there exist $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=H\circ Y$. Let $C$ be the projective curve in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by $h_{F,H}$, which is irreducible by assumption. Then $\phi=(X,Y):\Delta\dashrightarrow C$ is a non-constant rational map, where $\Delta$ is the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$. Since $\Delta\cong\mathbb{P}^{1}$ is non-singular and $C$ is projective, $\phi:\Delta\to C$ is a morphism. But $g(\Delta)=0$ and $g(C)>0$, so by the Riemann–Hurwitz formula, no non-constant morphism from $\Delta$ to $C$ exists, hence we get a contradiction. ∎
Theorem 3.6.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then the curve $h_{F}$ is irreducible and $g(h_{F})>0$. In particular, there exist no non-constant rational functions $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=F\circ Y$ in $\mathbb{C}(z)$, except when $X=Y$.
Proof.
By Proposition 3.2, ${\rm Mon}(F)\cong S_{m}$. In particular, ${\rm Mon}(F)$ is doubly transitive, so the curve $h_{F}$ is irreducible by [Pak11, Corollary 2.3]. By Fact 3.4, since $f$ is pre-simple, we have
$4-2g(h_{F})=m-(m-1-1)m^{2}+(m-1)((m-2)(m-1)+(m-2+2))=-m^{2}+5m-2,$
i.e., $g(h_{F})=(m-2)(m-3)/2$. For $m\geq 4$, we have $g(h_{F})>0$.
Suppose there exist $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ with $X\neq Y$ such that $F\circ X=F\circ Y$. Let $C$ be the irreducible projective curve in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by $h_{F}$. Since $X\neq Y$, the map $\phi=(X,Y):\Delta\dashrightarrow C$ is a non-constant rational map, where $\Delta$ is the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$. As before, $\phi:\Delta\to C$ is a morphism. But $g(\Delta)=0$ and $g(C)>0$, leading to a contradiction by the Riemann–Hurwitz formula. ∎
Self-compositions of a pre-simple polynomial behave well under decomposition.
Theorem 3.7.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Then for every $l\in\mathbb{Z}_{>0}$, every decomposition of $F^{\circ l}$ into a composition of indecomposable rational maps (of degree $\geq 2$) is equivalent to the decomposition $F^{\circ l}=F\circ\cdots\circ F$.
Proof.
The case $l=1$ follows from Proposition 3.2. Assume $l\geq 2$ and consider a decomposition $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ into indecomposable rational maps. By considering the totally ramified fixed point $\infty$, we may assume $F_{r},\dots,F_{1}$ are polynomials. By [ZM08, Theorem 1.3] and Proposition 3.2, we have $r=l$, $m=\deg(F_{l})=\cdots=\deg(F_{1})$, and ${\rm Mon}(F_{j})\cong S_{m}\cong{\rm Mon}(F)$ for all $1\leq j\leq l$. The tuple $(m=\deg(F_{l}),\dots,m=\deg(F_{1}))$ and the isomorphic type $S_{m}$ of ${\rm Mon}(F_{i})$ are invariants of complete decompositions of $F^{\circ l}$.
Suppose the decomposition $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ is not equivalent to $F^{\circ l}=F\circ\cdots\circ F$. Then by Ritt’s theorem on polynomial decomposition [FG22, Theorem 3.35] (see [Rit22, ZM08]), there exist $j\in\{1,\dots,l\}$ and linear polynomials $\sigma_{1}(z),\sigma_{2}(z)\in\mathbb{C}[z]$ such that $\sigma_{1}\circ F_{j}\circ\sigma_{2}(z)=z^{m}$. (Note that $\gcd(m,m)=m\neq 1$, so we need not consider the Chebyshev polynomial $T_{m}(z)$; see [FG22, §3.5.1 (M3)].) Then $S_{m}\cong{\rm Mon}(F_{i})\cong{\rm Mon}(\sigma_{1}\circ F_{j}\circ\sigma_{2})={\rm Mon}(z^{m})\cong\mathbb{Z}/m\mathbb{Z}$ (see [ZM08, Lemma 3.6]), contradicting $m\geq 3$. ∎
Remark 3.8.
In Theorem 3.7, we only require $m\geq 3$. The proof relies on Ritt’s theory on polynomial decompositions. For $m\geq 5$, an alternative proof following [Pak25a] is as follows:
We proceed by induction on $l$. The case $l=1$ follows from Proposition 3.2. Assume $l\geq 2$ and the result holds for l1l-1. Consider a decomposition $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ into indecomposable rational maps, with $F_{r},\dots,F_{1}$ polynomials.
Since $m\geq 5$, Theorem 3.5 implies that the algebraic curve $F(x)-F_{r}(y)=0$ is reducible (otherwise $F\circ X=F_{r}\circ Y$ has no solutions). By Proposition 3.2 and [Pak25a, Theorem 2.7], either $F_{r}=F\circ\mu$ for some $\mu\in{\rm PGL}_{2}(\mathbb{C})$, or $\deg(F_{r})=\binom{m}{k}$ for some 1<k<m11<k<m-1. If $F_{r}=F\circ\mu$, then $F^{\circ(l-1)}=(\mu^{-1}\circ F_{r-1})\circ\cdots F_{1}$ by Theorem 3.6, and the conclusion follows by induction. If $\deg(F_{r})=\binom{m}{k}$ for some 1<k<m11<k<m-1, then by [Pak25a, Theorem 2.9], since $m\geq 4$, there exists a prime $p$ dividing $\binom{m}{k}=\deg(F_{r})$ such that $p\nmid m$. But $p$ divides $m^{l}=\deg(F^{\circ l})$ since $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$, which is a contradiction. ∎
Corollary 3.9.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$, and let $G_{1},\dots,G_{r}\in\mathbb{C}(z)$ be rational maps of degree $\geq 2$ such that $F^{\circ l}=G_{r}\circ\cdots\circ G_{1}$ for some $l\in\mathbb{Z}_{>0}$. Then there exist $s_{1},\dots,s_{r}\in\mathbb{Z}_{>0}$ and Möbius transformations $\nu_{1},\dots,\nu_{r-1}\in\mathbb{C}(z)$ of degree $1$ such that
$G_{r}=F^{\circ s_{r}}\circ\nu_{r-1},\quad G_{j}=\nu_{j}^{-1}\circ F^{\circ s_{j}}\circ\nu_{j-1}\ (1<j<r),\quad G_{1}=\nu_{1}^{-1}\circ F^{\circ s_{1}}.$
The $s_{1},\dots,s_{r}$ are uniquely determined by $m=\deg(F)$ and the degrees of the $G_{j}$’s, with $\sum_{j=1}^{r}s_{j}=l$. If $G_{1},\dots,G_{r}$ are polynomials, then $\nu_{1},\dots,\nu_{r-1}$ can be chosen as linear polynomials.
Proof.
Decompose each $G_{i}$ ($1\leq i\leq r$) into indecomposable rational functions and apply Theorem 3.7. If $G_{1},\dots,G_{r}$ are polynomials, then using the totally ramified fixed point $\infty$, we see that $\nu_{1},\dots,\nu_{r-1}$ are linear polynomials. ∎
We now recall certain (semi)groups (under composition) associated with a rational map $F(z)\in\mathbb{C}(z)$ of degree $m\geq 2$, following [Pak25a].
Let
$C(F):=\{g\in\mathbb{C}(z)\setminus\mathbb{C}\mid F\circ g=g\circ F\}$
be the semigroup of non-constant rational maps commuting with $F$. Let
$C_{\infty}(F):=\bigcup_{l=1}^{\infty}C(F^{\circ l})$
be the semigroup of non-constant rational maps commuting with some iterate of $F$. Define
${\rm Aut}(F):=C(F)\cap{\rm PGL}_{2}(\mathbb{C})\quad\text{and}\quad{\rm Aut}_{\infty}(F):=C_{\infty}(F)\cap{\rm PGL}_{2}(\mathbb{C}).$
Both ${\rm Aut}(F)$ and ${\rm Aut}_{\infty}(F)$ are groups. Note that the semigroup $\left\langle{\rm Aut}_{\infty}(F),F\right\rangle$ generated by ${\rm Aut}_{\infty}(F)\cup\{F\}$ is contained in $C_{\infty}(F)$.
For $g\in\mathbb{C}(z)$ of degree $\geq 2$, let $\mu_{g}$ denote the measure of maximal entropy of $g$ (see [Lyu83, Mañ83]). Define
$E_{0}(F):=\{\sigma\in{\rm PGL}_{2}(\mathbb{C})\mid\sigma_{*}\mu_{F}=\mu_{F}\}$
and
$E(F):=E_{0}(F)\cup\{g\in\mathbb{C}(z)\mid\deg(g)\geq 2,\mu_{g}=\mu_{F}\}.$
Then $E_{0}(F)$ is a group, and $E(F)$ is a semigroup (see [CM21]).
Define $G(F)$ as the subgroup of ${\rm PGL}_{2}(\mathbb{C})$ given by
$\{\sigma\in{\rm PGL}_{2}(\mathbb{C}):\exists\tau\in{\rm PGL}_{2}(\mathbb{C}),\,F\circ\sigma=\tau\circ F\}.$
Define $G_{0}(F)$ as the maximal subgroup of ${\rm PGL}_{2}(\mathbb{C})$ such that for every $\sigma\in G_{0}(F)$, there exists $\tau\in G_{0}(F)$ with $F\circ\sigma=\tau\circ F$. Clearly, $G_{0}(F)\subseteq G(F)$.
Lemma 3.10.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then $G_{0}(F)$ is finite. Define $\gamma:G(F)\to{\rm PGL}_{2}(\mathbb{C})$ by $\sigma\mapsto\tau_{\sigma}$, where $\tau_{\sigma}$ is the unique Möbius transformation such that $F\circ\sigma=\tau_{\sigma}\circ F$. Then the restriction $\gamma:G_{0}(F)\to\gamma(G_{0}(F))$ is a group automorphism of $G_{0}(F)$.
Proof.
By [Pak20b, §4], the group $G(F)$ is finite of order bounded in terms of $m=\deg(F)$, unless $(\alpha\circ F\circ\beta)(z)=z^{m}$ for some $\alpha,\beta\in{\rm PGL}_{2}(\mathbb{C})$. Since $F$ is pre-simple, for any Möbius transformations $\alpha$ and $\beta$, the map $\alpha\circ F\circ\beta$ is pre-simple and has exactly $m$ critical values in $\widehat{\mathbb{C}}$. But $z^{m}$ has only $2(<m)$ critical values in $\widehat{\mathbb{C}}$. Hence, $G(F)$ is finite, and so is its subgroup $G_{0}(F)$. By definition, $\gamma(G_{0}(F))\subseteq G_{0}(F)$. By Theorem 3.6, $\gamma$ is injective for $m\geq 4$. Thus, $\gamma:G_{0}(F)\to G_{0}(F)$ is a bijection. It is easy to check that $\gamma$ is a group homomorphism, hence an automorphism. ∎
Corollary 3.11.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$, and let $s=\#G_{0}(F)\in\mathbb{Z}_{>0}$. Then $G_{0}(F)\subseteq{\rm Aut}(F^{\circ s})$.
Proof.
By Lemma 3.10, $\gamma:G_{0}(F)\to\gamma(G_{0}(F))$ is a group automorphism of the finite group $G_{0}(F)$ of order $s$, so $\gamma^{\circ s}={\rm Id}_{G_{0}(F)}$. For every $\sigma\in G_{0}(F)$, we have
$F^{\circ s}\circ\sigma=\gamma^{\circ s}(\sigma)\circ F^{\circ s}=\sigma\circ F^{\circ s},$
hence $G_{0}(F)\subseteq{\rm Aut}(F^{\circ s})$. ∎
By definition, it is easy to see that pre-simple polynomials of degree $\geq 4$ are non-exceptional.
Lemma 3.12.
Every pre-simple polynomial $F\in\mathbb{C}[z]$ of degree $m\geq 4$ is non-exceptional.
Proof.
Suppose $F$ is a pre-simple polynomial of degree $m\geq 4$ that is exceptional. Since $F$ is a polynomial, it is not a Lattès map. Then $F$ is of monomial type, i.e., conjugate to $z^{\pm m}$ or $\pm T_{m}(z)$, where $T_{m}$ is the Chebyshev polynomial of degree $m$. As $F$ is a polynomial, it cannot be conjugate to $z^{-m}$. Hence, $F$ is conjugate to $z^{m}$, $T_{m}(z)$, or $T_{m}(z)$ by a linear polynomial. Then $z^{m}$ or $T_{m}(z)$ must be pre-simple. But $z^{m}$ is not pre-simple for $m\geq 3$. The Chebyshev polynomial $T_{m}$ satisfies $(4-z^{2})F^{\prime}(z)^{2}=m^{2}(4-(F(z))^{2})$, so it has at most $2$ distinct critical values ($\pm 2$) in $\mathbb{C}$. Since 2<m12<m-1, $T_{m}$ is not pre-simple and we get a contradiction. ∎
Remark 3.13.
Every polynomial of degree $2$ is automatically pre-simple. Note that the Chebyshev polynomial $T_{3}(z)=z^{3}-3z$ of degree $3$ is pre-simple.
The following theorem describes the relations between the semigroups introduced above for a pre-simple polynomial of degree $\geq 4$, analogous to [Pak25a, Theorem 1.2].
Theorem 3.14.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then
$G_{0}(F)={\rm Aut}(F^{\circ s})={\rm Aut}_{\infty}(F)=E_{0}(F),\quad C_{\infty}(F)=\left\langle{\rm Aut}_{\infty}(F),F\right\rangle=E(F),$
where $s=\#G_{0}(F)\in\mathbb{Z}_{>0}$.
Proof.
We have $G_{0}(F)\subseteq{\rm Aut}(F^{\circ s})\subseteq{\rm Aut}_{\infty}(F)\subseteq E_{0}(F)$, where the first inclusion is by Corollary 3.11, the second by definition, and the third by [Pak25a, (31)].
To show $E_{0}(F)\subseteq G_{0}(F)$, let $\sigma\in E_{0}(F)$. Then $F\circ\sigma\in E(F)$. By Lemma 3.12, $F$ is non-exceptional. By [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ with $k_{1}=k_{2}+l$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ(F\circ\sigma)^{\circ l}$. Applying Theorem 3.6 recursively, we get $F^{\circ l}=F^{\circ(k_{1}-k_{2})}=(F\circ\sigma)^{\circ l}$. By Theorem 3.7, there exist $\mu_{1},\dots,\mu_{l-1}\in{\rm PGL}_{2}(\mathbb{C})$ such that
$F\circ\sigma=F\circ\mu_{l-1},\quad F\circ\sigma=\mu_{i}^{-1}\circ F\circ\mu_{i-1}\ (1<i<l),\quad F\circ\sigma=\mu_{1}^{-1}\circ F.$
Set $\mu_{1}(z)=z$ when $l=1$. Then $\nu:=\mu_{1}^{-1}$ satisfies $F\circ\sigma=\nu\circ F$. We claim $\nu\in E_{0}(F)$. For $l=1$, this is clear. For $l\geq 2$, note that $F^{\circ l}=(F\circ\sigma)^{\circ l}=(\nu\circ F)^{\circ l}$. By [Pak25a, Lemma 3.3], $\nu\in{\rm Aut}(F^{\circ l})\subseteq{\rm Aut}_{\infty}(F)\subseteq E_{0}(F)$. Hence, $E_{0}(F)\subseteq G_{0}(F)$.
Thus, $G_{0}(F)={\rm Aut}(F^{\circ s})={\rm Aut}_{\infty}(F)=E_{0}(F)$.
By [Pak25a, (32)], $C_{\infty}(F)\subseteq E(F)$. Note that
$C_{\infty}(F)\cap{\rm PGL}_{2}(\mathbb{C})={\rm Aut}_{\infty}(F)=E_{0}(F)=E(F)\cap{\rm PGL}_{2}(\mathbb{C}).$
To show $C_{\infty}(F)=E(F)$, it suffices to prove $E(F)\setminus E_{0}(F)\subseteq C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})$. Let $g\in E(F)\setminus E_{0}(F)$. Then $\deg(g)\geq 2$ and $\mu_{g}=\mu_{F}$. By [Lev90], there exist integers $k_{1},l>0$ and $k_{2}\geq 0$ such that $F^{\circ k_{1}}=F^{\circ k_{2}}\circ g^{\circ l}$. Clearly, $k_{1}>k_{2}$. Applying Theorem 3.6 recursively, we get $F^{\circ(k_{1}-k_{2})}=g^{\circ l}$. Then $g$ commutes with $F^{\circ(k_{1}-k_{2})}=g^{\circ l}$, so $g\in C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})$. Hence, $C_{\infty}(F)=E(F)$.
Clearly, $\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\subseteq C_{\infty}(F)$. For maps of degree $1$,
$\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\cap{\rm PGL}_{2}(\mathbb{C})={\rm Aut}_{\infty}(F)=C_{\infty}(F)\cap{\rm PGL}_{2}(\mathbb{C}).$
To show $\left\langle{\rm Aut}_{\infty}(F),F\right\rangle=C_{\infty}(F)$, it suffices to prove
$C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})\subseteq\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\setminus{\rm PGL}_{2}(\mathbb{C}).$
Let $G\in C_{\infty}(F)\setminus{\rm PGL}_{2}(\mathbb{C})$. By [Rit23], there exist integers $k,l>0$ such that $F^{\circ k}=G^{\circ l}$. By Corollary 3.9, there exists $\mu_{1}\in{\rm PGL}_{2}(\mathbb{C})$ such that $G=\mu_{1}^{-1}\circ F^{\circ s}$, where $s=k/l\in\mathbb{Z}_{>0}$ (take $\mu_{1}(z)=z$ and $s=k$ when $l=1$). Then
$F^{\circ sl}=F^{\circ k}=G^{\circ l}=(\mu_{1}^{-1}\circ F^{\circ s})^{\circ l}.$
By [Pak25a, Lemma 3.3], $\mu_{1}^{-1}\in{\rm Aut}(F^{\circ sl})\subseteq{\rm Aut}_{\infty}(F)$. Then $G=\mu_{1}^{-1}\circ F^{\circ s}\in\left\langle{\rm Aut}_{\infty}(F),F\right\rangle\setminus{\rm PGL}_{2}(\mathbb{C})$.
We have shown that $C_{\infty}(F)=\left\langle{\rm Aut}_{\infty}(F),F\right\rangle=E(F)$ holds. ∎
We now consider general polynomials. Here, “general” means the property holds for polynomials in a non-empty Zariski open subset of the parameter space ${\rm Poly}^{m}(\mathbb{C})$.
Lemma 3.15.
For every integer $m\geq 2$, a general polynomial $F\in\mathbb{C}[z]$ of degree $m$ is pre-simple.
Proof.
Up to conjugation by linear polynomials, we only need to consider monic centered polynomials of degree $m$, i.e., polynomials of the form
$F(z)=z^{m}+a_{m-2}z^{m-2}+\cdots+a_{0},$
where $a_{m-2},\dots,a_{0}\in\mathbb{C}$. The parameter space of monic centered polynomials of degree $m$ is
${\rm Poly}_{mc}^{m}=\{(a_{0},\dots,a_{m-2})\}\cong\mathbb{A}^{m-1}.$
The moduli space ${\rm MPoly}^{m}$ is the quotient of ${\rm Poly}_{mc}^{m}$ by the finite cyclic group $\mathbb{U}_{m-1}$ of (m1)(m-1)-th roots of unity acting diagonally on $\mathbb{A}^{m-1}\cong{\rm Poly}_{mc}^{m}$ by
$\lambda\cdot(a_{0},a_{1},\dots,a_{m-2})=(\lambda a_{0},a_{1},\dots,\lambda^{3-m}a_{m-2}).$
See [FG22, §2.1] for more details. Let $R(t)$ be the resultant ${\rm Res}_{z}(F^{\prime}(z),F(z)-t)$ in $z$. Then
$R(t)=m^{m}\prod_{\zeta:F^{\prime}(\zeta)=0}(F(\zeta)-t)\in(\mathbb{C}(a_{0},\dots,a_{m-2}))^{{\rm alg}}[t].$
Set $Z:={\rm Res}_{t}(R(t),R^{\prime}(t))\in\mathbb{C}[a_{0},\dots,a_{m-2}]\setminus\{0\}$ (the existence of pre-simple $f\in{\rm Poly}_{mc}^{m}(\mathbb{C})$ implies $Z\neq 0$). The resultant $Z$ defines a hypersurface in ${\rm Poly}_{mc}^{m}\cong\mathbb{A}^{m-1}$, corresponding to polynomials that are not pre-simple. Hence, a general polynomial of degree $m$ is pre-simple. ∎
Lemma 3.16.
For every integer $m\geq 3$, a general polynomial $F\in\mathbb{C}[z]$ of degree $m$ satisfies $G(F)=1$.
Proof.
Embed ${\rm Rat}_{m}$ into $\mathbb{P}^{2m+1}$ via the coefficients of rational maps (see [Sil12]). By the proof of [Pak25a, Lemma 3.7], there exists a closed subvariety $Z$ of $\mathbb{P}^{2m+1}$ over $\mathbb{C}$ such that
$\{F\in{\rm Rat}_{m}(\mathbb{C}):G(F)\neq 1\}=Z\cap{\rm Rat}_{m}(\mathbb{C})$
and $F\notin Z$ for some monic centered polynomial $F\in{\rm Poly}_{mc}^{m}(\mathbb{C})$ of degree $m$. The conclusion follows. ∎
Lemmas 3.15, 3.16, and Theorem 3.14 immediately imply:
Theorem 3.17.
For every integer $m\geq 4$, a general polynomial $F\in\mathbb{C}[z]$ of degree $m$ satisfies $E_{0}(F)={\rm Aut}_{\infty}(F)=G_{0}(F)=1$ and $E(F)=C_{\infty}(F)=\left\langle F\right\rangle$.
The following lemma generalizes Lemma 3.12:
Lemma 3.18.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then for every $r\in\mathbb{Z}_{>0}$, the polynomial $F^{\circ r}$ is not a generalized Lattès map.
Proof.
Suppose $F^{\circ r}$ is a generalized Lattès map for some $r\in\mathbb{Z}_{>0}$, and let $\mathcal{O}$ be an orbifold on $\widehat{\mathbb{C}}$ with a non-constant ramification function $\nu$ such that $F^{\circ r}:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map. Set $c^{*}(\mathcal{O})=c(\mathcal{O})\setminus\{\infty\}$. Since $\mathcal{O}$ is good, $c^{*}(\mathcal{O})$ is non-empty. Let $k:=\#c^{*}(\mathcal{O})\in\mathbb{Z}_{>0}$ and define
$\mathcal{A}:=\{z\in\mathbb{C}:F^{\circ r}(z)\in c^{*}(\mathcal{O}),\deg_{z}(F)=\cdots=\deg_{F^{\circ(r-1)}(z)}(F)=1\}.$
Since $F$ is pre-simple, induction shows $\mathcal{A}\geq(m-2)^{r}k\geq 2k$. As $f:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map, we have $\mathcal{A}\subseteq c^{*}(\mathcal{O})$. Then $2k\leq\#\mathcal{A}\leq\#c^{*}(\mathcal{O})=k$, which is a contradiction. Hence, $F^{\circ r}$ is not a generalized Lattès map for any $r\in\mathbb{Z}_{>0}$. ∎
Remark 3.19.
According to [Pak23, §2.3], for a rational map $f\in\mathbb{C}(z)$ of degree $\geq 2$ not of monomial type, $f$ is a generalized Lattès map if and only if $f^{\circ r}$ is a generalized Lattès map for some $r\in\mathbb{Z}_{>0}$, which is equivalent to $f^{\circ r}$ being a generalized Lattès map for all $r\in\mathbb{Z}_{>0}$. Thus, by Lemma 3.12, it suffices to prove Lemma 3.18 for $r=1$.
We now state a result on decomposing polynomials involving pre-simple ones.
Theorem 3.20.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$, $G\in\mathbb{C}[z]$ a polynomial of degree $\geq 4$, and $X\in\mathbb{C}[z]\setminus\mathbb{C}$ a non-constant polynomial such that $X\circ G=F^{\circ r}\circ X$ for some $r\in\mathbb{Z}_{>0}$. Then $l:=\log_{m}(\deg(X))$ is a non-negative integer, and there exists $\nu\in{\rm PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ l}\circ\nu$ and $G=\nu^{-1}\circ F^{\circ r}\circ\nu$. If $G$ and $X$ are polynomials, we may take $\nu$ to be a linear polynomial.
Proof.
If $\deg(X)=1$, the result is trivial. Assume $\deg(X)\geq 2$. By Lemma 3.18, $F^{\circ r}$ is not a generalized Lattès map. Applying [Pak25a, Theorem 4.1], there exist $Y\in\mathbb{C}(z)\setminus\mathbb{C}$ and $d\in\mathbb{Z}_{>0}$ such that $X\circ Y=F^{\circ rd}$. By Corollary 3.9, $l:=\log_{m}(\deg(X))$ is a positive integer, and there exists $\mu\in{\rm PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ l}\circ\mu$. Then
$F^{\circ(r+l)}\circ\mu=F^{\circ r}\circ X=X\circ G=F^{\circ l}\circ\mu\circ G.$
By Theorem 3.6, we get $F^{\circ r}\circ\mu=\mu\circ G$, so $G=\mu^{-1}\circ F^{\circ r}\circ\mu$. If $G$ and $X$ are polynomials, then $Y$ is also a polynomial, and by Corollary 3.9, we may take $\mu$ to be a linear polynomial. ∎
Lemma 3.21.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then for every $k\in\mathbb{Z}_{>0}$, the map $\gamma:{\rm Aut}(F^{\circ k})\to{\rm Aut}(F^{\circ k})$ (as in Lemma 3.10) is a group automorphism.
Proof.
By Theorem 3.14, ${\rm Aut}_{\infty}(F)=G_{0}(F)$. Fix $k\in\mathbb{Z}_{>0}$. Then ${\rm Aut}(F^{\circ k})\subseteq{\rm Aut}_{\infty}(F)=G_{0}(F)$. For every $\nu\in{\rm Aut}(F^{\circ k})$, we have
$F^{\circ k}\circ\gamma(\nu)\circ F$ $=F^{\circ k}\circ F\circ\nu=F\circ(F^{\circ k}\circ\nu)$
$=F\circ\nu\circ F^{\circ k}=\gamma(\nu)\circ F\circ F^{\circ k}=\gamma(\nu)\circ F^{\circ k}\circ F,$
so $F^{\circ k}\circ\gamma(\nu)=\gamma(\nu)\circ F^{\circ k}$. Hence, $\gamma(\nu)\in{\rm Aut}(F^{\circ k})$. By Lemma 3.10, the map $\gamma:{\rm Aut}(F^{\circ k})\to{\rm Aut}(F^{\circ k})$ is an injective group endomorphism on the finite group ${\rm Aut}(F^{\circ k})$, hence an automorphism. ∎
We now describe periodic curves for endomorphisms $(F_{1},F_{2})$ on $(\mathbb{P}^{1})^{2}$, where $F_{1},F_{2}\in\mathbb{C}[z]$ are pre-simple of the same degree $\geq 4$.
Theorem 3.22.
Let $F_{1},F_{2}\in\mathbb{C}[z]$ be pre-simple polynomials of the same degree $m\geq 4$, and let $C\subset(\mathbb{P}^{1})^{2}$ be an irreducible algebraic curve over $\mathbb{C}$ that is not a vertical or horizontal line. Then for every $k\in\mathbb{Z}_{>0}$, the following are equivalent:
  1. (i)
    $(F_{1},F_{2})^{\circ k}(C)=C$;
  2. (ii)
    There exist $s\in\mathbb{Z}_{\geq 0}$, $\alpha\in{\rm PGL}_{2}(\mathbb{C})$, and $\nu\in{\rm Aut}(F_{1}^{\circ k})$ such that F_2^k=αF_1^kα^-1, and $C$ is one of the graphs y=(ανF_1^s)(x) or x=(νF_1^sα^-1)(y), where $x$ and $y$ are the coordinates on $(\mathbb{P}^{1})^{2}$.
Proof.
(ii) $\Rightarrow$ (i) is trivial.
Assume (i) holds for some $k\in\mathbb{Z}_{>0}$. By Lemma 3.18, $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are not generalized Lattès maps. By [Pak25a, Theorem 4.2], there exist non-constant rational maps $X_{1},X_{2},Y_{1},Y_{2},B\in\mathbb{C}(z)\setminus\mathbb{C}$ and $\tilde{k}\in\mathbb{Z}_{\geq 0}$ such that:
  • $F_{j}^{\circ k}\circ X_{j}=X_{j}\circ B$, $B\circ Y_{j}=Y_{j}\circ F_{j}^{\circ k}$, $X_{j}\circ Y_{j}=F_{j}^{\circ k\tilde{k}}$, $Y_{j}\circ X_{j}=B^{\circ\tilde{k}}$ for $j=1,2$;
  • The map $t\mapsto(X_{1}(t),X_{2}(t))$ parametrizes $C$.
By Theorem 3.20, there exist $\beta,\alpha\in{\rm PGL}_{2}(\mathbb{C})$ and $d_{1},d_{2}\in\mathbb{Z}_{\geq 0}$ such that
$X_{1}=F_{1}^{\circ d_{1}}\circ\beta,\quad X_{2}=F^{\circ d_{2}}\circ\alpha,\quad\text{and}\quad\beta^{-1}\circ F_{1}^{\circ k}\circ\beta=B=\alpha^{-1}\circ F_{2}^{\circ k}\circ\alpha.$
Conjugating $B$ suitably, we may assume $\beta(z)=z$. Then $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),(F_{2}^{\circ d_{2}}\circ\alpha)(t))$ and $F_{1}^{\circ k}=B=\alpha^{-1}\circ F_{2}^{\circ k}\circ\alpha$. Since
$\mu_{\alpha^{-1}\circ F_{2}\circ\alpha}=\mu_{(\alpha^{-1}\circ F_{2}\circ\alpha)^{\circ k}}=\mu_{F_{1}^{\circ k}}=\mu_{F_{1}},$
we have $\alpha^{-1}\circ F_{2}\circ\alpha\in E(F_{1})=\left\langle{\rm Aut}_{\infty}(F_{1}),F_{1}\right\rangle$ by Theorem 3.14. Also, ${\rm Aut}_{\infty}(F_{1})=G_{0}(F)$ by Theorem 3.14. Hence, $\alpha^{-1}\circ F_{2}\circ\alpha=\tau\circ F_{1}$ for some $\tau\in{\rm Aut}_{\infty}(F_{1})=G_{0}(F)$. Since $F_{1}^{\circ k}=(\alpha^{-1}\circ F_{2}\circ\alpha)^{\circ k}=(\tau\circ F_{1})^{\circ k}$, we get $\tau\in{\rm Aut}(F_{1}^{\circ k})$ by [Pak25a, Lemma 3.3]. Note that $F_{2}=\alpha\circ\tau\circ F_{1}\circ\alpha^{-1}$ and $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),\alpha\circ(\tau\circ F_{1})^{\circ d_{2}}(t))$. By Lemma 3.21, $(\tau\circ F_{1})^{\circ d_{2}}=\tau^{\prime}\circ F_{1}^{\circ d_{2}}$ for some $\tau^{\prime}\in{\rm Aut}(F_{1}^{\circ k})$. Then $C$ is parametrized by $t\mapsto(F_{1}^{\circ d_{1}}(t),(\alpha\circ\tau^{\prime}\circ F_{1}^{\circ d_{2}}(t))$.
If $d_{1}\leq d_{2}$, then $C$ is parametrized by $t\mapsto(t,(\alpha\circ\tau^{\prime}\circ F_{1}^{\circ(d_{2}-d_{1})}(t))$, and we are done by setting $\nu:=\tau^{\prime}$ and $s:=d_{2}-d_{1}$.
If $d_{1}>d_{2}$, then by Lemma 3.21,
$F_{1}^{\circ(d_{1}-d_{2})}\circ(\tau^{\prime})^{-1}\circ\alpha^{-1}=\tau^{\prime\prime}\circ F_{1}^{\circ(d_{2}-d_{1})}\circ\alpha^{-1}$
for some $\tau^{\prime\prime}\in{\rm Aut}(F_{1}^{\circ k})$. Then $C$ is parametrized by
$t\mapsto((\tau^{\prime\prime}\circ F_{1}^{\circ(d_{1}-d_{2})}\circ\alpha^{-1})(t),t),$
and we are done by setting $\nu:=\tau^{\prime\prime}$ and $s:=d_{1}-d_{2}$. ∎
Following [FG22], two polynomials $f,g\in\mathbb{C}[z]$ of degree $\geq 2$ are intertwined if there exists an algebraic curve $Z\subset(\mathbb{A}^{1})^{2}$ (possibly reducible) whose projections to both axes are onto, and $Z$ is invariant under the endomorphism $(f,g):(\mathbb{A}^{1})^{2}\to(\mathbb{A}^{1})^{2}$. The following theorem is a polynomial version of [JX25, Theorem 3.3]:
Theorem 3.23 (Theorem 1.5).
For every integer $d\geq 4$, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for every $f\in U$:
  1. (1)
    $f$ is pre-simple;
  2. (2)
    for every pre-simple $g\in{\rm Poly}^{d}(\mathbb{C})$, if $f$ and $g$ are intertwined, then $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Proof.
By Theorem 3.14, Lemma 3.15, and Lemma 3.16, there exists a non-empty Zariski open subset $W_{d}$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for every $f\in W_{d}$, the polynomial $f$ is pre-simple and
(3.1) $\{g\in{\rm Rat}_{d}(\mathbb{C}):\mu_{g}=\mu_{f}\}=\{f\}.$
We show that $W_{d}$ also satisfies (2). Let $f\in W_{d}$ and $g\in{\rm Poly}^{d}(\mathbb{C})$ be a pre-simple polynomial intertwined with $f$. By Theorem 3.22, there exist $k\in\mathbb{Z}_{>0}$ and $\alpha\in{\rm PGL}_{2}(\mathbb{C})$ such that
$f^{\circ k}=\alpha\circ g^{\circ k}\circ\alpha^{-1}=(\alpha\circ g\circ\alpha^{-1})^{\circ k}.$
Then
$\mu_{f}=\mu_{f^{\circ k}}=\mu_{(\alpha\circ g\circ\alpha^{-1})^{\circ k}}=\mu_{\alpha\circ g\circ\alpha^{-1}}.$
By (3.1), $f=\alpha\circ g\circ\alpha^{-1}$. Since $\alpha$ is a degree-$1$ polynomial, $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Hence, $W_{d}$ has the desired property. ∎
3.2. The case of degree $2\leq d\leq 3$
Having analyzed general polynomials of degree $d\geq 4$ in §3.1 using pre-simplicity, we now consider general polynomials of degree $2\leq d\leq 3$, largely motivated by Pakovich’s work [Pak25b] on general rational maps of degree $2\leq d\leq 3$. We show that for $d\in\{2,3\}$, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$, the polynomials $f$ and $g$ are intertwined if and only if they are conjugate. This result is also a direct corollary of [FG22, Theorems 3.51 and 3.52] (see [GNY19, Thoerem 1.4]).
For $d=2$, the moduli space ${\rm MPoly}^{2}\cong\mathbb{A}^{1}$ has dimension $1$, and every $f\in{\rm Poly}^{2}(\mathbb{C})$ is conjugate to a unique polynomial of the form $\phi_{c}(z)=z^{2}+c$ with $c\in\mathbb{C}$. Thus, degree $2$ polynomials are easier to study due to the simple normal forms $\phi_{c}$. For $d=3$, Theorem 3.7 holds, which concerns decompositions of iterates of a general polynomial and relies on Ritt’s theory. For both $d=2$ and $d=3$, we show that such a Zariski open subset $U$ exists.
For a non-constant rational map $f\in\mathbb{C}(z)$, define the groups
$\Sigma(f):=\{\sigma\in{\rm PGL}_{2}(\mathbb{C}):f\circ\sigma=f\}\quad\text{and}\quad\Sigma_{\infty}(f):=\bigcup_{k\geq 1}\Sigma(f^{\circ k}).$
Clearly, $\Sigma(f)\subseteq\Sigma_{\infty}(f)$, and this inclusion is preserved under conjugation: for $\tau\in{\rm PGL}_{2}(\mathbb{C})$,
$\Sigma(\tau\circ f\circ\tau^{-1})=\tau\circ\Sigma(f)\circ\tau^{-1}\subseteq\tau\circ\Sigma_{\infty}(f)\circ\tau^{-1}=\Sigma_{\infty}(\tau\circ f\circ\tau^{-1}).$
Lemma 3.24.
For every $c\in\mathbb{C}^{\times}$, we have $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})=\{z,-z\}$, the cyclic group of order two generated by z-z.
Proof.
Clearly, $\{z,-z\}\subseteq\Sigma(\phi_{c})\subseteq\Sigma_{\infty}(\phi_{c})$.
Since $c\neq 0$, the map $\phi_{c}$ has a unique totally ramified fixed point at $\infty$. Thus, every $\sigma\in\Sigma(\phi_{c})$ fixes $\infty$, so $\sigma$ is a linear polynomial. Write $\sigma(z)=az+b$ with $a\in\mathbb{C}^{\times}$, $b\in\mathbb{C}$. The equality $\phi_{c}\circ\sigma=\phi_{c}$ implies $\sigma(z)=z$ or $\sigma(z)=-z$. Hence, $\{z,-z\}=\Sigma(\phi_{c})$.
We now show by induction on $n\in\mathbb{Z}_{>0}$ that $\Sigma(\phi_{c})=\Sigma(\phi_{c}^{\circ n})$ for all $n$, so $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})$. The case $n=1$ is trivial. Assume $n\geq 2$ and $\Sigma(\phi_{c})=\Sigma(\phi_{c}^{\circ(n-1)})$. Clearly, $\Sigma(\phi_{c})\subseteq\Sigma(\phi_{c}^{\circ n})$. To show $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$, let $\sigma\in\Sigma(\phi_{c}^{\circ n})$. Since $c\neq 0$, $\phi_{c}^{\circ n}$ has a unique totally ramified fixed point at $\infty$, so $\sigma$ fixes $\infty$ and is a linear polynomial: $\sigma(z)=az+b$ with $a\in\mathbb{C}^{\times}$, $b\in\mathbb{C}$. By [Pak25b, Lemma 3.4] and $\phi_{c}\circ\phi_{c}^{\circ(n-1)}=\phi_{c}\circ(\phi_{c}^{\circ(n-1)}\circ\sigma)$, we have either $\phi_{c}^{\circ(n-1)}=\phi_{c}^{\circ(n-1)}\circ\sigma$ or $\phi_{c}^{\circ(n-1)}=-\phi_{c}^{\circ(n-1)}\circ\sigma$. If $\phi_{c}^{\circ(n-1)}=\phi_{c}^{\circ(n-1)}\circ\sigma$, then by induction, $\sigma\in\Sigma(\phi_{c}^{\circ(n-1)})=\Sigma(\phi_{c})$, so $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$.
Suppose $\phi_{c}^{\circ(n-1)}=-\phi_{c}^{\circ(n-1)}\circ\sigma$. Comparing coefficients:
  • The coefficient of $z^{2^{n-1}}$ gives $a^{2^{n-1}}=-1$.
  • The coefficient of $z^{2^{n-1}-1}$ gives $b=0$.
  • The coefficient of $z^{2^{n-1}-2}$ gives $2^{n-2}c=-a^{2^{n-1}-2}2^{n-2}c$, so $a^{2^{n-1}-2}=-1$ (since $c\neq 0$).
Then
$1=a^{2^{n-1}}=(a^{2})^{2^{n-2}}=(a^{2^{n-1}}/a^{2^{n-1}-2})^{2^{n-2}}=(-1/-1)^{2^{n-2}}=1,$
which is a contradiction. Hence, $\Sigma(\phi_{c}^{\circ n})\subseteq\Sigma(\phi_{c})$. By induction, we obtain $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})$. ∎
The following two theorems show that for $2\leq d\leq 3$, a general polynomial $F\in\mathbb{C}[z]$ of degree $d$ satisfies certain “good” properties.
Theorem 3.25.
There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{2}(\mathbb{C})$ such that for every $F\in U$:
  1. (1)
    $\Sigma(F)=\Sigma_{\infty}(F)$ is cyclic of order $2$;
  2. (2)
    if $G(z)\in\mathbb{C}(z)$ is a rational map of degree $2$ such that $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$, then $G=F$;
  3. (3)
    for every $l\in\mathbb{Z}_{>0}$, every decomposition of $F^{\circ l}$ into indecomposable rational maps (of degree $\geq 2$) is equivalent to $F^{\circ l}=F\circ\cdots\circ F$.
Proof.
Let $U:=\{a_{1}^{2}-2a_{1}\neq 4a_{0}a_{2}\}$, a non-empty Zariski open subset of
${\rm Poly}^{2}(\mathbb{C})=\{a_{0}z^{2}+a_{1}z+a_{2}\}\cong\{(a_{0},a_{1},a_{2}):a_{0}\neq 0\}.$
A simple computation shows that $a_{0}z^{2}+a_{1}z+a_{2}\in{\rm Poly}^{2}(\mathbb{C})$ is conjugate to some $z^{2}+c$ with $c\neq 0$ if and only if it lies in $U$.
Let $F(z)=a_{0}z^{2}+a_{1}z+a_{2}\in U$, and let $c=c(F)\in\mathbb{C}^{\times}$ be such that $F$ is conjugate to $\phi_{c}$. Choose $\tau\in{\rm PGL}_{2}(\mathbb{C})$ with $\tau\circ F\circ\tau^{-1}=\phi_{c}$.
(1) By Lemma 3.24, $\Sigma(\phi_{c})=\Sigma_{\infty}(\phi_{c})=\{z,-z\}$ is cyclic of order $2$, so $\Sigma(F)=\Sigma_{\infty}(F)$ is also cyclic of order $2$.
[…]
$\phi_{c}^{\circ k}\circ\mu_{\tilde{G}}=\tilde{G}^{\circ k}\circ\mu_{\tilde{G}}=\tilde{G}^{\circ k}=\phi_{c}^{\circ k},$
so $\mu_{\tilde{G}}\in\Sigma_{\infty}(\phi_{c})=\{\pm z\}$ by Lemma 3.24. Since $\mu_{\tilde{G}}(z)\neq z$, we have $\mu_{\tilde{G}}(z)=-z$. By [Pak25b, Lemma 3.5], $\tilde{G}=\nu\circ\phi_{c}$ for some $\nu\in{\rm PGL}_{2}$. From $\phi_{c}^{\circ k}=\tilde{G}^{\circ k}=(\nu\circ\phi_{c})^{\circ k}$, we get $\phi_{c}^{\circ(k-1)}=(\nu\circ\phi_{c})^{\circ(k-1)}\circ\nu$, so
$\phi_{c}^{\circ k}=\phi_{c}\circ\phi_{c}^{\circ(k-1)}=\phi_{c}\circ(\nu\circ\phi_{c})^{\circ(k-1)}\circ\nu=(\phi_{c}\circ\nu)^{\circ k}.$
The same argument shows that the generator of $\Sigma(\phi_{c}\circ\nu)$ is $\mu_{\phi_{c}\circ\nu}(z)=-z$. Note that
$\phi_{c}\circ\nu\circ(\nu^{-1}\circ(-z)\circ\nu)=\phi_{c}\circ(-z)\circ\nu=\phi_{c}\circ\nu,$
so $\nu^{-1}\circ(-z)\circ\nu(z)=\mu_{\phi_{c}\circ\nu}(z)=-z$, hence $\nu(-z)=-\nu(z)$. Thus, $\nu(z)=\lambda z$ or $\nu(z)=\lambda/z$ for some $\lambda\in\mathbb{C}^{\times}$. If $\nu(z)=\lambda/z$, then $\infty$ is not a totally ramified fixed point of $(\lambda/(z^{2}+c))^{\circ k}=\phi_{c}^{\circ k}$ (since $c\neq 0$), contradicting that $\phi_{c}^{\circ k}$ is a polynomial. So $\nu(z)=\lambda z$. Then
(3.2) $(z^{2}+c)^{\circ k}=\phi_{c}^{\circ k}(z)=\tilde{G}^{\circ k}(z)=(\nu\circ\phi_{c})^{\circ k}(z)=(\lambda z^{2}+\lambda c)^{\circ k}.$
Comparing the coefficient of $z^{2^{k}}$ of the two sides of (3.2) gives $\lambda^{2^{k}-1}=1$. If $k=1$, then $\lambda=1$. If $k\geq 2$, the coefficient of $z^{2^{k}-4}$ gives
$2^{k-3}(2^{k}-2)c^{2}+2^{k-2}c=2^{k-3}(2^{k}-2)\lambda^{2^{k}-1}c^{2}+2^{k-2}\lambda^{2^{k}-3}c,$
so $\lambda^{2}=1$ (since $c\neq 0$ and $\lambda^{2^{k}-1}=1$). Then $\lambda=(\lambda^{2})^{2^{k-1}}/\lambda^{2^{k}-1}=1$. Hence, $\nu(z)=z$, so $\tilde{G}=\phi_{c}$ and $G=F$.
(3) We proceed by induction on $l$. For $l=1$, the result is trivial since $\deg(F^{\circ l})=2$ is prime. Assume $l\geq 2$ and the result holds for l1l-1. Let $F^{\circ l}=F_{r}\circ\cdots\circ F_{1}$ be a decomposition into indecomposable rational maps. It suffices to show $F_{1}=\mu\circ F$ for some $\mu\in{\rm PGL}_{2}(\mathbb{C})$. By [Pak25b, Corollary 2.2], $\deg(F_{1})=2$. Let $\mu_{F_{1}}$ be the generator of $\Sigma(F_{1})$ (cyclic of order $2$ by [Pak25b, Lemma 3.4 (i)]). Then
$F^{\circ l}\circ\mu_{F_{1}}=F_{r}\circ\cdots\circ F_{1}\circ\mu=F_{r}\circ\cdots\circ F_{1}=F^{\circ l},$
so ${\rm Id}\neq\mu_{F_{1}}\in\Sigma_{\infty}(F)=\Sigma(F)$. By [Pak25b, Lemma 3.5], $F_{1}=\mu\circ F$. The result follows by induction. ∎
Theorem 3.26.
There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{3}(\mathbb{C})$ such that for every $F\in U$:
  1. (1)
    $F(z)$ is pre-simple and $\Sigma(F)=\Sigma_{\infty}(F)=1$;
  2. (2)
    if $G(z)\in\mathbb{C}(z)$ is a rational map of degree $3$ such that $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$, then $G=F$;
  3. (3)
    for every $l\in\mathbb{Z}_{>0}$, every decomposition of $F^{\circ l}$ into indecomposable rational maps (of degree $\geq 2$) is equivalent to $F^{\circ l}=F\circ\cdots\circ F$.
Proof.
By Lemmas 3.15 and 3.16, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{3}(\mathbb{C})$ such that every $F\in U$ is pre-simple and $G(F)=1$. Then (3) follows from Theorem 3.7.
Let $F\in U$. We prove (1) and (2).
(1) $F$ is pre-simple by construction. After conjugation, assume $F(z)=z^{3}+az+b$ with $a,b\in\mathbb{C}$. Since $F$ is pre-simple, $F^{\prime}(z)=3z^{3}+a$ has no multiple roots, so $a\neq 0$. Let $n\geq 1$ and $\sigma\in\Sigma(\phi_{c}^{\circ n})$. Then $\sigma$ is a linear polynomial (as in the proof of Lemma 3.24). Write $\sigma(z)=ez+f$ with $e\in\mathbb{C}^{\times}$, $f\in\mathbb{C}$. By induction, the leading terms of $(z^{3}+az+b)^{\circ k}$ are:
(3.3) $(z^{3}+az+b)^{\circ k}=z^{3^{k}}+0\cdot z^{3^{k}-1}+3^{k-1}az^{3^{k}-2}+O(z^{3^{k}-3}).$
Substituting into $(z^{3}+az+b)^{\circ k}\circ(ez+f)=(z^{3}+az+b)^{\circ k}$ and comparing coefficients of $z^{3^{k}},z^{3^{k}-1},z^{3^{k}-2}$ shows $(e,f)=(1,0)$, so $\sigma(z)=z$. Hence, $\Sigma(F)=\Sigma_{\infty}(F)=1$.
(2) Let $G\in\mathbb{C}(z)$ be a rational map of degree $3$ with $G^{\circ k}=F^{\circ k}$ for some $k\in\mathbb{Z}_{>0}$. Since $\deg(F)=\deg(G)=3$ is prime, both are indecomposable. By (3), the decomposition $F^{\circ k}=G\circ\cdots\circ G$ implies $G=\nu\circ F$ for some $\nu\in{\rm PGL}_{2}(\mathbb{C})$. Then $F^{\circ k}=(\nu\circ F)^{\circ k}$, so $F^{\circ(k-1)}=(\nu\circ F)^{\circ(k-1)}\circ\nu$. Composing with $F$ gives $F^{\circ k}=(F\circ\nu)^{\circ k}$. Applying (3) again, $F\circ\nu=\delta\circ F$ for some $\delta\in{\rm PGL}_{2}(\mathbb{C})$. Then $\nu\in G(F)=1$, so $\nu(z)=z$ and $G=F$. ∎
Lemma 3.27.
For every integer $d\geq 2$, a general polynomial $F\in\mathbb{C}[z]$ of degree $d$ is not a generalized Lattès map.
Proof.
For $d\geq 4$, the result follows from Lemmas 3.18 and 3.15.
If $d=2$, then $F$ is a generalized Lattès map if and only if it is exceptional (Proposition 2.4). Exceptional polynomials of degree $2$ are conjugate to $z^{2}$ or $T_{2}(z)=z^{2}-2$. Since ${\rm MPoly}^{2}(\mathbb{C})\cong\mathbb{C}$ has dimension $1>0$, a general $F$ is not exceptional.
If $d=3$, then $F$ is exceptional if and only if it is conjugate to $z^{3}$ or $\pm T_{3}(z)=\pm(z^{3}-3z)$. Thus a general $F$ is not exceptional. By Proposition 2.4 and $\deg(F)=3$, a non-exceptional $F$ is a generalized Lattès map if and only if it is conjugate to $zR(z)^{2}=z(az+b)^{2}$ for some $a,b\in\mathbb{C}^{\times}$ (since $F$ is not conjugate to $z^{3}$). Thus, $F$ is a generalized Lattès map if and only if there exist $a,b,e\in\mathbb{C}^{\times}$ and $f\in\mathbb{C}$ such that
(3.4) $F(z)=\frac{(ez+f)(a(ez+f)+b)^{2}-f}{e}.$
Setting $\lambda=ae\in\mathbb{C}^{\times}$ and $\mu=af\in\mathbb{C}$, this becomes
(3.5) $F(z)=\lambda^{2}z^{3}+\lambda(3\mu+2b)z^{2}+(\mu+b)(3\mu+b)z+\frac{\mu}{\lambda}(\mu^{2}+2\mu b+b^{2}-1).$
There are only $3$ parameters $\lambda,b\in\mathbb{C}^{\times}$ and $\mu\in\mathbb{C}$, while ${\rm Poly}^{3}(\mathbb{C})=\{a_{0}z^{3}+a_{1}z^{2}+a_{2}z+a_{3}:(a_{0},a_{1},a_{2},a_{3})\in\mathbb{C}^{\times}\times\mathbb{C}^{3}\}$ has dimension $4>3$. Hence, a general $F$ is not a generalized Lattès map. ∎
Theorem 3.28.
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for every $F\in U$, if $B,X\in\mathbb{C}(z)\setminus\mathbb{C}$ and $k\in\mathbb{Z}_{>0}$ satisfy $F^{\circ k}\circ X=X\circ B$, then there exist $m\in\mathbb{Z}_{\geq 0}$ and $\mu\in{\rm PGL}_{2}(\mathbb{C})$ such that $X=F^{\circ m}\circ\mu$ and $B=\mu^{-1}\circ F^{\circ k}\circ\mu$.
Proof.
By Lemma 3.27 and Theorems 3.25 and 3.26, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that every $F\in U$ is not a generalized Lattès map and satisfies conditions (1)–(3) of Theorem 3.25 (for $d=2$) or Theorem 3.26 (for $d=3$).
Let $F\in U$, and suppose $B,X\in\mathbb{C}(z)\setminus\mathbb{C}$ satisfy $F^{\circ k}\circ X=X\circ B$ for some $k\in\mathbb{Z}_{>0}$. By Remark 3.19, $F^{\circ k}$ is not a generalized Lattès map. Applying [Pak23, Proposition 3.3], there exist $Y\in\mathbb{C}(z)\setminus\mathbb{C}$ and $r\in\mathbb{Z}_{\geq 0}$ such that $X\circ Y=F^{\circ kr}$. By (3) of Theorem 3.25 or 3.26, $X=F^{\circ m}\circ\delta$ for some $m\in\mathbb{Z}_{\geq 0}$ and $\delta\in{\rm PGL}_{2}(\mathbb{C})$. Then
(3.6) $F^{\circ(k+m)}=F^{\circ m}\circ\delta\circ B\circ\delta^{-1}.$
Again by (3), there exists $\nu\in{\rm PGL}_{2}(\mathbb{C})$ such that $\delta\circ B\circ\delta^{-1}=\nu^{-1}\circ F^{\circ k}$. Then $F^{\circ(k+m)}=F^{\circ m}\circ\nu^{-1}\circ F^{\circ k}$, so $F^{\circ m}=F^{\circ m}\circ\nu$ and $\nu\in\Sigma_{\infty}(F)$ (if $m=0$, then $\nu(z)=z\in\Sigma_{\infty}(F)$). By (1), $\nu\in\Sigma(F)$. Let $\mu:=\nu\circ\delta$. Then
$X=F^{\circ m}\circ\delta=F^{\circ m}\circ\nu\circ\delta=F^{\circ m}\circ\mu,$
and
$B=\delta^{-1}\circ\nu^{-1}\circ F^{\circ k}\circ\delta=\delta^{-1}\circ\nu^{-1}\circ F^{\circ k}\circ\nu\circ\delta=\mu^{-1}\circ F^{\circ k}\circ\mu.$
Theorem 3.29.
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(F_{1},F_{2})\in U\times U$ and every irreducible algebraic curve $C\subset\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ that is not a horizontal or vertical line, the following are equivalent:
  1. (1)
    $C$ is $(F_{1},F_{2})$-periodic;
  2. (2)
    there exist $\alpha\in{\rm PGL}_{2}(\mathbb{C})$ and $s\in\mathbb{Z}_{\geq 0}$ such that $F_{2}=\alpha\circ F_{1}\circ\alpha^{-1}$ and $C$ is one of the graphs y=(αF_1^s)(x) or x=(F_1^sα^-1)(y), where $x$ and $y$ are the coordinates on $\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$.
In particular, every $(F_{1},F_{2})$-periodic irreducible curve is $(F_{1},F_{2})$-invariant.
Proof.
By Lemma 3.27 and Theorems 3.25 and 3.26, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that every $F\in U$ is not a generalized Lattès map and satisfies (1)–(3) of Theorem 3.25 (for $d=2$) or 3.26 (for $d=3$). By the proof of Theorem 3.28, the conclusion of that theorem holds for every $F\in U$.
Let $F_{1},F_{2}\in U$ have the same degree $d$, and let $C\subset\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ be an irreducible curve that is not a vertical or horizontal line.
Assume (1): $C$ is $(F_{1},F_{2})$-periodic. Then for some $k\in\mathbb{Z}_{>0}$, $C$ is $(F_{1}^{\circ k},F_{2}^{\circ k})$-invariant. By Remark 3.19, both $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are not a generalized Lattès map. Applying [Pak23, Theorem 1.1], there exist non-constant rational maps $X_{1},X_{2}\in\mathbb{C}(z)$ such that:
  • $F_{i}^{\circ k}\circ X_{i}=X_{i}\circ B$ for $i=1,2$;
  • $t\mapsto(X_{1}(t),X_{2}(t))$ parametrizes $C$.
By Theorem 3.28, there exist $\alpha_{1},\alpha_{2}\in{\rm PGL}_{2}(\mathbb{C})$ and $m_{1},m_{2}\in\mathbb{Z}_{\geq 0}$ such that
$X_{1}=F_{1}^{\circ m_{1}}\circ\alpha_{1},\quad X_{2}=F_{2}^{\circ m_{2}}\circ\alpha_{2},\quad\alpha_{1}^{-1}\circ F_{1}^{\circ k}\circ\alpha_{1}=B=\alpha_{2}^{-1}\circ F_{2}^{\circ k}\circ\alpha_{2}.$
Let $\alpha=\alpha_{2}\circ\alpha_{1}^{-1}$. Then
$F_{2}^{\circ k}=\alpha\circ F_{1}^{\circ k}\circ\alpha^{-1}=(\alpha\circ F_{1}\circ\alpha^{-1})^{\circ k},$
so $F_{2}=\alpha\circ F_{1}\circ\alpha^{-1}$ by (2) of Theorem 3.25 or 3.26. The curve $C$ is parametrized by
$t\mapsto(X_{1}(t),X_{2}(t))=(F_{1}^{\circ m_{1}}\circ\alpha_{1}(t),\alpha\circ F_{1}^{\circ m_{2}}\circ\alpha_{1}(t)),$
so $t\mapsto(F_{1}^{\circ m_{1}}(t),\alpha\circ F_{1}^{\circ m_{2}}(t))$ also parametrizes $C$.
If $m_{1}\geq m_{2}$, set $s=m_{1}-m_{2}\in\mathbb{Z}_{\geq 0}$; then $t\mapsto(F_{1}^{\circ s}\circ\alpha^{-1}(t),t)$ parametrizes $C$. If $m_{1}<m_{2}$, set $s=m_{2}-m_{1}\in\mathbb{Z}_{>0}$; then $t\mapsto(t,\alpha\circ F_{1}^{\circ s}(t))$ parametrizes $C$. Thus, $C$ is one of the stated graphs.
Conversely, if (2) holds, then $C$ is clearly $(F_{1},F_{2})$-invariant. ∎
Remark 3.30.
Compared to Theorem 3.22, the form of periodic curves in Theorem 3.29 is simpler due to Theorems 3.25(2) and 3.26(2). By (2) of Theorem 3.25 or 3.26, for two general polynomials $F_{1},F_{2}\in\mathbb{C}[z]$ of the same degree $d\in\{2,3\}$, $F_{1}$ and $F_{2}$ are conjugate if and only if $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are conjugate for some $k\in\mathbb{Z}_{>0}$. If $F$ satisfies (2) of Theorem 3.25 or 3.26, then ${\rm Aut}(F)={\rm Aut}_{\infty}(F)$. For $\nu\in{\rm Aut}(F)$, we have $\nu\circ F\circ\nu^{-1}=F$, so we may replace $\alpha$ in Theorem 3.22 by $\alpha\circ\nu$ or $\alpha\circ\nu^{-1}$ when $\nu\in{\rm Aut}(F_{1})$.
Proposition 3.31 (Proposition 1.6).
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$ with $f$ and $g$ intertwined, we have $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Proof.
Take $U$ as in Theorem 3.29. The result follows directly. ∎
3.3. Proof of Theorem 1.4
Using Theorem 3.23 or Proposition 3.31 in place of [FG22, Theorems 3.51 and 3.52], we obtain a proof of Theorem 1.4 following Ji-Xie [JX25].
Proof of Theorem 1.4.
Let $d\geq 2$ be an integer. Assume by contradiction that $\tilde{\tau}_{d}$ is not generically injective. By Theorem 3.23 or Proposition 3.31, there exists a non-empty Zariski open subset $U$ of ${\rm MPoly}^{d}(\mathbb{C})$ such that for all $f,g\in{\rm Poly}^{d}(\mathbb{C})$ with $f],[g]\in U$,
[…]
Take a non-empty Zariski open subset $W$ of the Zariski closure of $\tilde{\tau}_{d}(U)$ such that $\tilde{\tau}_{d}^{-1}(W)\subseteq U$ and $\tilde{\tau}_{d}:\tilde{\tau}_{d}^{-1}(W)\to W$ is a finite étale morphism of degree at least $2$. After shrinking $U$, we may assume $U=\tilde{\tau}_{d}^{-1}(W)$.
It is well known that the PCF locus $\{f\in{\rm Poly}^{d}(\mathbb{C}):f\text{ is }{\rm PCF}\}$ is Zariski-dense in ${\rm Poly}^{d}$. As in [JX25, Proof of Theorem 1.3], we can find two non-isotrivial algebraic families $h_{1},h_{2}$ of degree $d$ polynomials parameterized by an irreducible affine curve $C$ with $h_{1}(C)\cup h_{2}(C)\subseteq U$ and
$\tilde{\tau}_{d}\circ h_{1}=\tilde{\tau}_{d}\circ h_{2}:C\to\mathbb{A}_{\mathbb{C}}^{N}$
such that for every $t\in C$, $h_{1}(t)\neq h_{2}(t)$ in ${\rm MPoly}^{d}$, and the PCF locus
$\{t\in C:h_{1}(t)\text{ is }{\rm PCF}\}$
for $h_{1}$ is Zariski-dense in $C$. Here $N=N_{d,1}+\cdots N_{d,m_{d}}$.
By [JXZ25, Theorem 1.14], whether a rational map is PCF is determined by its length spectrum (hence its multiplier spectrum). So
$\{t\in C:h_{2}(t)\text{ is }{\rm PCF}\}=\{t\in C:h_{1}(t)\text{ is }{\rm PCF}\},$
which is Zariski-dense in $C$. Applying the variant of DAO (Dynamical André-Oort) conjecture for curves proved in [JX25, Theorem 3.4], the set
$\{t\in C:h_{1}(t)\text{ and }h_{2}(t)\text{ are intertwined}\}$
is infinite. (The condition that $h_{1},h_{2}$ are not families of flexible Lattès maps holds automatically for polynomials.) Then by (3.7), the set
$\{t\in C:h_{1}(t)=h_{2}(t)\text{ in }{\rm MPoly}^{d}\}$
is infinite, contradicting the construction of $h_{1}$ and $h_{2}$. ∎
3.4. A case of the Zariski-dense orbit conjecture
As a byproduct, we now present a proof of the Zariski-dense orbit conjecture (ZDO) for a general split polynomial endomorphism on $(\mathbb{P}^{1})^{2}$ with all factors of the same degree $d\geq 2$, which is essentially due to Pakovich [Pak25a, Pak25b].
Conjecture 3.32 (ZDO).
Let $k$ be an algebraically closed field of characteristic 0. Given an irreducible quasi-projective variety $X$ over $k$ and a dominant rational self-map $f$ on $X$. If $\{g\in k(X):g\circ f=g\}=k$, where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
Theorem 3.33.
Fix an integer $d\geq 2$. For a general pair of polynomials $F_{1},F_{2}\in{\rm Poly}^{d}(\mathbb{C})$, the endomorphism $(F_{1},F_{2}):\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}\to\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ has no irreducible periodic curves other than vertical or horizontal lines. In particular, for every $x\in\mathbb{P}^{1}(\mathbb{C})\setminus{\rm PrePer}(F_{1})$ and $y\in\mathbb{P}^{1}(\mathbb{C})\setminus{\rm PrePer}(F_{2})$, the Zariski closure of the forward orbit $\mathcal{O}_{(F_{1},F_{2})}((x,y))$ is $\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$.
Proof.
For $d\in\{2,3\}$, the result follows from Theorem 3.29.
Assume $d\geq 4$. By Lemmas 3.15 and 3.16, a general $f\in{\rm Poly}^{d}(\mathbb{C})$ is pre-simple with $G(f)=1$. Let $F_{1},F_{2}$ be pre-simple polynomials of degree $m$ with $G(F_{1})=1$, $G(F_{2})=1$, and $F_{1}]\neq[F_{2}]$. Such pairs are general in ${\rm Poly}^{d}(\mathbb{C})\times{\rm Poly}^{d}(\mathbb{C})$.
Suppose $C$ is an irreducible $(F_{1},F_{2})$-periodic curve in $\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ that is not a vertical or horizontal line. By Theorem 3.22, $F_{2}^{\circ k}=\alpha\circ F_{1}^{\circ k}\circ\alpha^{-1}$ for some $k\in\mathbb{Z}_{>0}$ and $\alpha\in{\rm PGL}_{2}(\mathbb{C})$. By Theorem 3.14 and $G_{0}(F_{2})\subseteq G(F_{2})=1$, we have $E(F_{2})=\left\langle F_{2}\right\rangle$. Since
$\mu_{\alpha\circ F_{1}\circ\alpha^{-1}}=\mu_{(\alpha\circ F_{1}\circ\alpha^{-1})^{\circ k}}=\mu_{F_{2}^{\circ k}}=\mu_{F_{2}},$
we have $\alpha\circ F_{1}\circ\alpha^{-1}\in E(F_{2})=\left\langle F_{2}\right\rangle$. Comparing degrees, $\alpha\circ F_{1}\circ\alpha^{-1}=F_{2}$, contradicting $F_{1}]\neq[F_{2}]$. ∎
Remark 3.34.
Xie [Xie17] proved ZDO for all dominant polynomial endomorphisms $f:\mathbb{A}_{k}^{2}\to\mathbb{A}_{k}^{2}$ over an algebraically closed field $k$ of characteristic 0 using valuative techniques.
4. Polynomials with the same multiplier spectrum
4.1. Marked critical points and a theorem of Favre-Gauthier
On the parameter space ${\rm Poly}^{d}$ (or the moduli space ${\rm MPoly}^{d}$), the critical points do not form global morphisms. This causes difficulties when working with information about critical points (e.g., PCF maps and stable preperiodic critical points). To address this, it is necessary to consider critically marked polynomials via a base change of the parameter space (or the moduli space). In this subsection, we gather classical constructions of critically marked polynomials; see [FG22, §2.1].
A critically marked polynomial of degree $d\geq 2$ is a tuple $(P,c_{0},\dots,c_{d-2})$ where $P\in{\rm Poly}^{d}$ and $c_{0},\dots,c_{d-2}$ are the critical points (other than $\infty$) of $P$, counted with multiplicity. Let ${\rm MPcrit}^{d}$ denote the quotient of the space of critically marked polynomials of degree $d$ by the natural action of the group $\rm Aff$} of affine transformations. There is a canonical forgetful morphism ${\rm MPcrit}^{d}\to{\rm MPoly}^{d}$.
Define the morphism
$\eta:\mathbb{A}^{d-1}\to{\rm MPcrit}^{d},(c,a)=(c_{1},\dots,c_{d-2},a)\mapsto[P_{c,a}(z)],$
where
$P_{c,a}(z)=\frac{1}{d}z^{d}+\sum_{j=2}^{d-1}(-1)^{d-j}\sigma_{d-j}(c)\frac{z^{j}}{j}+a^{d}$
and $\sigma_{j}(c)$ is the $j$-th elementary symmetric polynomial in $(c_{1},\dots,c_{d-2})$ for $1\leq j\leq d-2$. Note that the (finite) critical points of $P_{c,a}$ are exactly $c_{0}:=0,c_{1},\dots,c_{d-2}$. The morphism $\eta:\mathbb{A}^{d-1}\to{\rm MPcrit}^{d}$ is defined over $\mathbb{Q}$ and is d(d1)d(d-1)-to-one, so $\mathbb{A}^{d-1}$ is a finite ramified cover of ${\rm MPcrit}^{d}$.
The following theorem is a part of [FG22, Theorem B] and will be needed to prove Theorem 1.9.
Theorem 4.1 (Favre-Gauthier).
Let $(P,a)$ and $(Q,b)$ be non-exceptional active dynamical pairs parametrized by an irreducible algebraic curve $C$, of respective degrees $d,\delta\geq 2$, defined over a number field $K$. Then the following are equivalent:
  1. (1)
    The set $\{t\in C(\overline{K}):a(t)\in{\rm PrePer}(P(t))\text{ and }b(t)\in{\rm PrePer}(Q(t))\}$ is infinite;
  2. (2)
    There exist integers $N,M\geq 1$, $r,s\geq 0$, and families $R,\pi,\tau$ of polynomials of degree $\geq 1$ parametrized by $C$ such that τP^N=Rτ and πQ^M=Rπ with $\tau(P^{\circ N}(a))=\pi(Q^{\circ M}(b))$.
Here, a dynamical pair parametrized by $C$ of degree $d$ is a pair $(P,a)$ where $P:C\to{\rm MPoly}^{d}$ is a family of polynomials of degree $d$ over $C$, and $a:C\to\mathbb{A}^{1}$ is a morphism (i.e., a marked point on the family $P$). The dynamical pair $(P,a)$ is called active (or not stably preperiodic) if for all distinct integers $n,m\geq 0$ we have $(P(t))^{\circ n}(a(t))\neq(P(t))^{\circ m}(a(t))$ on $C$. It is called non-exceptional if the polynomial $P(t)$ is not exceptional for every $t\in U(\overline{K})$, where $U$ is a suitable dense open subset of $C$. If two pairs $(P,a)$ and $(Q,b)$ satisfy the property in (1), then they are called entangled.
4.2. Description of the non-injective locus
We now prove Theorem 1.9, relying on results of Pakovich and Theorem 4.1 of Favre-Gauthier.
From the proof of [JX25, Lemma 3.5], we obtain the following elementary lemma:
Lemma 4.2.
Let $f$ be a polynomial map of degree $d\geq 2$. Then its multiplier spectrum $(S_{n}(f))_{n=1}^{\infty}$ determines the number ${\rm sac}(f)\in\{0,1,\dots,2d-2\}$ of superattracting cycles of $f$ in $\mathbb{C}$ (counted without multiplicity). (The bound ${\rm sac}(f)\leq 2d-2$ follows from the Riemann–Hurwitz formula.)
Proposition 4.3.
Let $\phi:C\to{\rm MPoly}^{d}$ be a non-isotrivial family of polynomials of degree $d\geq 2$, parametrized by an irreducible curve $C$. Assume that $\phi$ factors through $\mathbb{A}^{d-1}\xrightarrow{\eta}{\rm MPcrit}^{d}\to{\rm MPoly}^{d}$. Let $m(\phi)\geq 0$ be the minimal nonnegative integer such that the subset $\{t\in C(\mathbb{C}):{\rm sac}(\phi(t))=m\}$ is infinite. Define
$Z(\phi):=\{t\in C(\mathbb{C}):{\rm sac}(\phi(t))>m(\phi)\},$
and let $Z^{\prime}(\phi)$ be the set of all $t\in C(\mathbb{C})$ such that $c_{j}$ is not a stably preperiodic marked point of $\phi$ but $c_{j}(t)\in{\rm Per}(\phi(t))$ for some $0\leq j\leq d-1$.
Then $m(\phi)$ and $Z(\phi)$ depend only on the multiplier spectrum $(S_{n}(\phi(t)))_{n\geq 1,t\in C(\mathbb{C})}$. Moreover, $Z(\phi)$ is an infinite subset of $C(\mathbb{C})$ and the symmetric difference
$Z(\phi)\Delta Z^{\prime}(\phi)=(Z(\phi)\cup Z^{\prime}(\phi))\setminus(Z(\phi)\cap Z^{\prime}(\phi))$
is finite.
Proof.
By Lemma 4.2, the number $m(t)$ depends only on the multiplier spectrum, and hence so does $Z(\phi)$. Note that $m(\phi)$ does not change after removing finitely many points of $C(\mathbb{C})$. In the following argument, we are free to remove finitely many points of $C(\mathbb{C})$.
Let $0\leq i,j\leq d-1$ be indices such that both $c_{i}$ and $c_{j}$ are stably periodic for $\phi$. Assume that $n,m\in\mathbb{Z}_{>0}$ are minimal such that $\phi(t)^{\circ n}(c_{i}(t))=c_{i}(t)$ and $\phi(t)^{\circ m}(c_{j}(t))=c_{j}(t)$ on all of $C$. For every integer $0\leq r\leq n$, the subset
$\{t\in C(\mathbb{C}):c_{j}(t)=\phi(t)^{\circ r}(c_{i}(t))\}$
is either finite or all of $C(\mathbb{C})$. After removing finitely many points of $C(\mathbb{C})$, we may assume that for all such pairs $(i,j)$, either
$\mathcal{O}_{\phi(t)}(c_{i}(t))=\mathcal{O}_{\phi(t)}(c_{j}(t))\quad\text{for all }t\in C(t),$
or
$\mathcal{O}_{\phi(t)}(c_{i}(t))\cap\mathcal{O}_{\phi(t)}(c_{j}(t))=\emptyset\quad\text{for all }t\in C(t).$
In the first case, write $i\sim j$. Let
$P(\phi):=\{0\leq i\leq d-1:\ c_{i}\text{ is stably periodic for }\phi\}.$
Then $\sim$ is an equivalence relation on $P(\phi)$.
Let $0\leq i\leq d-1$ be such that $c_{i}$ is stably preperiodic but not stably periodic for $\phi$. Assume $k>l\geq 0$ are minimal such that $\phi(t)^{\circ k}(c_{i}(t))=\phi(t)^{\circ l}(c_{i}(t))$ on all of $C$. Then $l>0$. For every $l\leq r\leq k$, the subset
$\{t\in C(\mathbb{C}):c_{i}(t)=\phi(t)^{\circ r}(c_{i}(t))\}$
is finite, since $c_{i}$ is not stably periodic. After removing finitely many points of $C(\mathbb{C})$, we may assume that
(4.1) $c_{i}(t)\notin{\rm Per}(\phi(t))\quad\text{for all }t\in C(\mathbb{C}).$
It is well known that PCF parameters do not form families in ${\rm MPoly}^{d}$. Thus,
(4.2) $A(\phi):=\{0\leq i\leq d-1:\ c_{i}\text{ is not stably preperiodic for }\phi\}\neq\emptyset.$
Fix any $i\in A(\phi)$. Similarly (after removing finitely many points of $C(\mathbb{C})$), we may assume that $c_{i}(t)\notin\mathcal{O}_{\phi(t)}(c_{k}(t))$ for all $t\in C(\mathbb{C})$ and all $k\in P(\phi)$. By [JX25, Lemma 2.4],
(4.3) $\{t\in C(\mathbb{C}):c_{i}(t)\in{\rm Per}(\phi(t))\}\ \text{is infinite.}$
On the other hand, since $c_{i}$ is not stably preperiodic for $\phi$, the set
$\{t\in C(\mathbb{C}):\phi(t)^{\circ n}(c_{i}(t))=c_{i}(t)\}$
is finite for every $n\geq 1$. Thus, by the uncountability of $C(\mathbb{C})$,
(4.4) $\bigcap_{k\in A(\phi)}\{t\in C(\mathbb{C}):c_{k}(t)\notin{\rm Per}(\phi(t))\}=+\infty.$
From the case-by-case analysis above, it is easy to see that
$m(\phi)=\#(P(\phi)/\sim).$
Indeed, we have ${\rm sac}(\phi(t))\geq\#(P(\phi)/\sim)$ for all $t\in C(\mathbb{C})$ by the analysis of stably periodic $c_{i}$’s, hence $m(\phi)\geq\#(P(\phi)/\sim)$. From (4.1) and (4.4), there are uncountably many $t\in C(\mathbb{C})$ with ${\rm sac}(\phi(t))=\#(P(\phi)/\sim)$. Hence $m(\phi)=\#(P(\phi)/\sim)$.
By (4.2) and (4.3), $Z^{\prime}(\phi)$ is infinite. For every $t\in Z^{\prime}(\phi)$,
${\rm sac}(\phi(t))\geq\#(P(\phi)/\sim)+1>m(\phi),$
so $Z^{\prime}(\phi)\subseteq Z(\phi)$; hence $Z(\phi)$ is also infinite. Conversely, for every $t\in C(\mathbb{C})\setminus Z^{\prime}(\phi)$, by (4.1) we have ${\rm sac}(\phi(t))=\#(P(\phi)/\sim)=m(\phi)$, so $Z(\phi)\subseteq Z^{\prime}(\phi)$. Therefore $Z(\phi)=Z^{\prime}(\phi)$, completing the proof. ∎
Proof of Theorem 1.9.
After a suitable base change of $C$ (restricting to a suitable non-empty affine open subset), we may assume that $\phi_{1}(t)$ and $\phi_{2}(t)$ are non-exceptional such that $(\phi_{1}(t),\phi_{2}(t))\in{\rm PNI}_{d}$ for every $t\in C(\mathbb{C})$, and that $\phi_{1},\phi_{2}$ factor through
$\mathbb{A}^{d-1}\xrightarrow{\eta}{\rm MPcrit}^{d}\to{\rm MPoly}^{d}.$
We continue to denote by $\phi_{1},\phi_{2}$ the corresponding morphisms $C\to\mathbb{A}^{d-1}$.
By Proposition 4.3, we have $m(\phi_{1})=m(\phi_{2})$, $Z(\phi_{1})=Z(\phi_{2})$, and the subsets $Z^{\prime}(\phi_{1})$ and $Z^{\prime}(\phi_{2})$ are infinite with finite symmetric difference $Z^{\prime}(\phi_{1})\Delta Z^{\prime}(\phi_{2})$. For $i=1,2$ and $k\in A(\phi_{i})$, set
$Z_{ik}=\{t\in C(\mathbb{C}):c_{k}(t)\in{\rm Per}(\phi_{i}(t))\}.$
Note that $Z_{ik}=\{t\in C(\overline{\mathbb{Q}}):c_{k}(t)\in{\rm Per}(\phi_{i}(t))\}$ since all the objects are defined over $\overline{\mathbb{Q}}$. Then $Z^{\prime}(\phi_{i})=\bigcup_{k\in A(\phi_{i})}Z_{ik}$ for $i=1,2$. Since $Z^{\prime}(\phi_{1})\cap Z^{\prime}(\phi_{2})$ is infinite, we can choose $k\in A(\phi_{1})$ and $l\in A(\phi_{2})$ such that $(Z_{1k}\cap Z_{2l})=+\infty$. After renumbering, assume $k=l=0$. Then 0 is not stably preperiodic for $\phi_{1}$ and $\phi_{2}$, and we obtain two active dynamical pairs $(\phi_{1},0)$ and $(\phi_{2},0)$ with
$\{t\in C(\mathbb{C}):0\in{\rm Per}(\phi_{1}(t))\cap{\rm Per}(\phi_{2}(t))\}=+\infty$
(since $(Z_{10}\cap Z_{20})=+\infty$). In the sense of Favre–Gauthier [FG22], the pairs $(\phi_{1},0)$ and $(\phi_{2},0)$ are entangled.
Applying Theorem 4.1, there exist $N,M\in\mathbb{Z}_{>0}$ and families $R,\tau,\pi$ of polynomials (of degree $\geq 1$) parametrized by $C$ such that
$\tau\circ\phi_{1}^{\circ N}=R\circ\tau,\qquad\pi\circ\phi_{2}^{\circ M}=R\circ\pi,\qquad\tau(\phi_{1}^{\circ N}(0))=\pi(\phi_{2}^{\circ M}(0)).$
Note that $d^{N}=\deg(\phi_{1}^{\circ N})=\deg(R)=\deg(\phi_{2}^{\circ M})=d^{M}$, hence $N=M$. By composing with $\phi_{1},\phi_{2}$, respectively, we may assume that the families $\tau$ and $\pi$ are of degree at least $2$.
Fix $t\in C(\mathbb{C})$. Write $h_{1}=\phi_{1}(t)^{\circ N}$ and $h_{2}=\phi_{2}(t)^{\circ N}$. For simplicity, we write $R,\tau,\pi$ for $R(t),\tau(t),\pi(t)$, respectively. Then
$\tau\circ h_{1}=R\circ\tau,\quad\pi\circ h_{2}=R\circ\pi,\quad\tau(h_{1}(0))=\pi(h_{2}(0)).$
Assume $h_{1}$ and $h_{2}$ are not equivalent. We will show that there exist positive integers $k_{1},k_{2},l$, a non-constant polynomial $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$, and $\zeta\in\mathbb{C}$, with
$2\leq k:={\rm lcm}(k_{1},k_{2})\leq d^{N},\quad\gcd(d,k)=1,\quad V(0)\neq 0,\quad\zeta^{k_{2}}=1,$
such that
$h_{1}\sim z^{l}\cdot V(z^{k_{1}})^{k_{1}^{\prime}}\quad\text{and}\quad h_{2}\sim\zeta\,z^{l}\cdot V(z^{k_{2}})^{k_{2}^{\prime}},$
where $k_{j}^{\prime}=k/k_{j}$ for $j=1,2$.
Since $h_{1}\not\sim h_{2}$, there exists $j\in\{1,2\}$ such that $h_{j}\not\sim R$. Without loss of generality, assume $h_{1}\not\sim R$. By Theorem 2.10, there exist non-constant rational maps $\tau_{0},\tau_{1},\tilde{h}_{1}$ satisfying:
  • $R$ is a generalized Lattès map;
  • $\tau=\tau_{0}\circ\tau_{1}$ with $\deg(\tau_{0})\geq 2$;
  • $\tilde{h}_{1}\circ\tau_{1}=\tau_{1}\circ h_{1}$, $R\circ\tau_{0}=\tau_{0}\circ\tilde{h}_{1}$, and $h_{1}\sim\tilde{h}_{1}$;
  • $(f=R,p=\tau_{0},g=\tau_{0},q=\tilde{h}_{1})$ is a good solution of $f\circ p=g\circ q$;
  • $R:\mathcal{O}_{2}^{\tau_{0}}\to\mathcal{O}_{2}^{\tau_{0}}$ and $\tilde{h}_{1}:\mathcal{O}_{1}^{\tau_{0}}\to\mathcal{O}_{1}^{\tau_{0}}$ are minimal holomorphic maps between orbifolds;
  • there exists $s\in\mathbb{Z}_{>0}$ such that $\tau_{1}$ is a compositional right factor of $h_{1}^{\circ s}$ and a compositional left factor of $\tilde{h}_{1}^{\circ s}$.
By conjugating by Möbius transformations, we may assume that $\tau_{1},\tau_{0}$, and $\tilde{h}_{1}$ are polynomials. Set $k_{1}=\deg(\tau_{0})\geq 2$.
We claim that, after further conjugation by a Möbius transformation, the polynomial $\tau_{0}$ has the form $\tau_{0}(z)=z^{k_{1}}$. If $R$ is exceptional, then by [FG22, Theorem 3.39], $h_{1}=\phi_{1}(t)^{\circ N}$ and $\phi_{1}(t)$ are also exceptional, contradicting our assumption. Thus $R$ is a non-exceptional polynomial which is a generalized Lattès map. Since $\tau_{0}:\mathcal{O}_{1}^{\tau_{0}}\to\mathcal{O}_{2}^{\tau_{0}}$ is a covering map between orbifolds, the Riemann–Hurwitz formula gives
$\chi(\mathcal{O}_{2}^{\tau_{0}})=\frac{\chi(\mathcal{O}_{1}^{\tau_{0}})}{k_{1}}<2$
because $k_{1}\geq 2$. In particular, the ramification function of $\mathcal{O}_{2}^{\tau_{0}}$ is not identically $1$. By Lemma 2.4, we have $\nu(\mathcal{O}_{2}^{\tau_{0}})=\{n,n\}$ for some integer $n\geq 2$. By definition of $\mathcal{O}_{2}^{\tau_{0}}$ and considering the point $\infty$, we see $k_{1}=\deg(\tau_{0})\in\nu(\mathcal{O}_{2}^{\tau_{0}})$, hence $n=k_{1}$. Since $\nu(\mathcal{O}_{2}^{\tau_{0}})=\{k_{1},k_{1}\}$, the polynomial $\tau_{0}$ has exactly two critical values in $\widehat{\mathbb{C}}$, hence is of the form $\tau_{0}(z)=az^{k_{1}}+c$, where $a\in\mathbb{C}^{\times}$ and $c\in\mathbb{C}$. Replacing $(\tau_{0},R,\pi)$ by $(L\circ\tau_{0},L\circ R\circ L^{-1},L\circ\pi)$, where L(z)=(zc)/aL(z)=(z-c)/a, we may assume $\tau_{0}(z)=z^{k_{1}}$.
By Theorem 2.8 and Remark 2.9, there exists $R_{1}(z)\in\mathbb{C}[z]$ such that
$R(z)=z^{r_{1}}\cdot R_{1}(z)^{k_{1}}\quad\text{and}\quad\tilde{h}_{1}(z)=z^{r_{1}}\cdot R_{1}(z^{k_{1}}),$
for some integer $r_{1}\geq 1$ with $\gcd(r_{1},k_{1})=1$.
Similarly, if $h_{2}\not\sim R$, there exist $R_{2}(z),\tilde{h}_{2}\in\mathbb{C}[z]$ such that
$h_{2}\sim\tilde{h}_{2},\quad R(z)=z^{r_{2}}\cdot R_{2}(z)^{k_{2}},\quad\tilde{h}_{2}(z)=z^{r_{2}}\cdot R_{2}(z^{k_{2}}),$
for some integers $r_{2}\geq 1$ and $k_{2}\geq 2$ with $\gcd(r_{2},k_{2})=1$. If $h_{2}\sim R$, set $\tilde{h}_{2}=R_{2}=R$, $r_{2}=0$, and $k_{2}=1$.
Set $k={\rm lcm}(k_{1},k_{2})\geq k_{1}\geq 2$ and $k_{j}^{\prime}=k/k_{j}$ for $j=1,2$. Then
(4.5) $\tilde{h}_{1}(z^{k_{1}^{\prime}})^{k_{1}}=R(z^{k})=\tilde{h}_{2}(z^{k_{2}^{\prime}})^{k_{2}}.$
Write $R_{j}(z)=z^{l_{j}}V_{j}(z)$ with $l_{j}\in\mathbb{Z}_{\geq 0}$ and $V_{j}(z)\in\mathbb{C}[z]$ satisfying $V_{j}(0)\neq 0$, for $j=1,2$. Then (4.5) becomes
$z^{(r_{1}+k_{1}l_{1})k}\cdot V_{1}(z^{k}))^{k_{1}}=z^{(r_{2}+k_{2}l_{2})k}\cdot V_{2}(z^{k})^{k_{2}}.$
Comparing orders at 0, we get $r_{1}+k_{1}l_{1}=r_{2}+k_{2}l_{2}=:l$. Hence $V_{1}(z^{k})^{k_{1}}=V_{2}(z^{k})^{k_{2}}$, so $V_{1}(z)^{k_{1}}=V_{2}(z)^{k_{2}}$. Examining irreducible factors of $V_{1}(z)$, there exists $V(z)\in\mathbb{C}[z]$ such that $V_{1}(z)=V(z)^{k_{1}^{\prime}}$. Then $V_{2}(z)^{k_{2}}=V(z)^{k}$, which implies $V_{2}(z)=\zeta\cdot V(z)^{k_{2}^{\prime}}$ for some $k_{2}$-th root of unity $\zeta\in\mathbb{C}^{\times}$. Therefore
$\tilde{h}_{1}(z)=z^{l}\cdot V(z^{k_{1}})^{k_{1}^{\prime}}\quad\text{and}\quad\tilde{h}_{2}(z)=\zeta\,z^{l}\cdot V(z^{k_{2}})^{k_{2}^{\prime}}.$
Note that $\gcd(r_{1},k_{1})=1$ implies $\gcd(d,k_{1})=1$, since
$d^{N}=\deg(R)=r_{1}+k_{1}\cdot\deg(R_{1}).$
Similarly, $\gcd(d,k_{2})=1$. Hence $\gcd(d,k_{1}k_{2})=\gcd(d,k)=1$. If $V(z)$ is a constant $c\in\mathbb{C}^{\times}$, then $l=d^{N}\geq 2$ and $h_{1}(z)\sim c^{k_{1}^{\prime}}z^{l}\sim z^{l}$, so all multipliers of $h_{1}$ are in $\mathbb{Z}$. By [JXZ25, Theorem 1.5], $h_{1}=\phi_{1}(t)^{\circ N}$ is exceptional and so is $\phi_{1}(t)$, contradicting our assumption. Thus $V(z)$ is non-constant. In particular,
$d^{N}=\deg(h_{1})=l+k\deg(V)\geq k\geq 2.$
After a suitable iterate, we can make further reductions. Set $V_{1}(z)=V(z)$. For $n\geq 1$, define $V_{n+1}(z)=V(z)^{l^{n}}\cdot V_{n}(z^{l}V(z)^{k})$ inductively. Then $V_{n}(z)\in\mathbb{C}[z]\setminus\mathbb{C}$ and $V_{n}(0)\neq 0$ for all $n\geq 1$. By induction, it is clear that for every $n\geq 1$,
(4.6) $\tilde{h}_{1}^{\circ n}(z)=z^{l^{n}}\cdot V_{n}(z^{k_{1}})^{k_{1}^{\prime}}\quad\text{and}\quad\tilde{h}_{2}^{\circ n}(z)=\zeta^{1+\cdots+l^{n-1}}\cdot z^{l^{n}}\cdot V_{n}(z^{k_{2}})^{k_{2}^{\prime}}.$
Since $\gcd(d,k)=1$ and $l=d^{N}-k\deg(V)$, $\gcd(l,k)=1$. By Euler’s theorem, for the integer $n=k_{2}\varphi(k)\geq 1$, we have
$l^{n}\equiv 1\pmod{k}\quad\text{and}\quad 1+\cdots+l^{n-1}\equiv 0\pmod{k_{2}}$.
Let $n_{0}\geq 1$ be the minimal integer such that
$l^{n_{0}}\equiv 1\pmod{k}\quad\text{and}\quad 1+\cdots+l^{n_{0}-1}\equiv 0\pmod{k_{2}}$.
Replacing $(h_{i},\tilde{h}_{i},N,l,V,\zeta)$ by $(h_{i}^{\circ n_{0}},\tilde{h}_{i}^{\circ n_{0}},n_{0}N,l^{n_{0}},V_{n_{0}},1)$ ($i=1,2$), we may assume $l\equiv 1\pmod{k}$ and $\zeta=1$. We still assume $h_{1}\not\sim h_{2}$ and hence $\tilde{h}_{1}\not\sim\tilde{h}_{2}$. In particular, $k_{1}\neq k_{2}$. This concludes the proof. ∎
Remark 4.4.
Pakovich [Pak17, Theorem 1.6] (see also [FG22, Theorem 3.44]) showed that for every polynomial $P$ of degree $d\geq 2$, there exist a polynomial $P_{\rm min}$ of degree $d$ and a non-constant polynomial $\pi_{\rm min}$ with $P\geq_{\pi_{\rm min}}P_{\rm min}$ satisfying the following universal property: for any polynomial $Q\leq P$, there exist polynomials $\pi,\omega$ such that $P\geq_{\omega}Q\geq_{\pi}P_{\rm min}$ and $\pi_{\rm min}=\pi\circ\omega$. Furthermore, $\deg(\pi_{\rm min})$ has an upper bound depending only on $d$.
From the proof above, we immediately see that in Theorem 1.9 (2), the polynomials $(\phi_{1}(t)^{\circ N})_{\rm min}$ and $(\phi_{2}(t)^{\circ N})_{\rm min}$ are mutually semi-conjugate, hence
$(\phi_{1}(t)^{\circ N})_{\rm min}\sim(\phi_{2}(t)^{\circ N})_{\rm min}$
by [Pak17, Theorem 1.5].
Remark 4.5.
By [FG22, Theorem B], the positive integer $N$ in (1) or (2) has an upper bound depending on $\mathbb{Q}(\phi_{1},\phi_{2})$ and $d$, as $n_{0}\leq k_{2}\varphi(k)$.
Proof of Theorem 1.8.
Considering a non-empty affine subset of $C$, by Theorem 1.9, there exists a finite subset $S\subset C(\mathbb{C})$ such that for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})\setminus S$, $\phi_{1}(t):=f_{t}$ and $\phi_{2}(t):=g_{t}$ satisfy (1) or (2) of Theorem 1.9. By [FG22, 4. and 5. of Theorem 3.39, 3. of Proposition 3.41], $f_{t}$ and $g_{t}$ are intertwined if they satisfy Theorem 1.9 (1). Note that for all $Q\in\mathbb{C}[z]\setminus\{0\}$ and $m,n\geq 0$, the polynomials $z^{m}Q(z^{n})$ and $z^{m}Q(z)^{n}$ are intertwined [FG22, Theorem 3.39]. Hence by [FG22, Theorem 3.39], we conclude that $f_{t}$ and $g_{t}$ are intertwined if they satisfy Theorem 1.9 (2). Consequently, for all but finitely many $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined. Ji and Xie [JX25, Remark 1.6] showed that the intertwined locus is Zariski-closed in $\mathcal{M}_{d}\times\mathcal{M}_{d}$, hence also Zariski-closed in ${\rm MPoly}^{d}\times{\rm MPoly}^{d}$. Therefore, for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined. ∎
4.3. Ritt move
Case 2 of Theorem 1.9 involves polynomials arising from the “Ritt move” in the Ritt theory of polynomial decompositions [Rit22]; see [ZM08] and [FG22, §3.5.1]. Up to composing with linear polynomials, the only solutions of the equation $P\circ Q=\hat{P}\circ\hat{Q}$ in indecomposable $P,Q,\hat{P},\hat{Q}\in\mathbb{C}[z]$ (of degree $\geq 2$) are the trivial one $P\circ Q=P\circ Q$ and the nontrivial ones has the forms:
(4.7) $z^{n}\circ z^{s}R(z^{n})$ $=z^{s}R(z)^{n}\circ z^{n},$
(4.8) $T_{p}\circ T_{q}$ $=T_{q}\circ T_{p},$
[…]
With additional hypotheses, we can still deduce that the multiplier spectra coincide on an (infinite) arithmetic progression in case (2) of Theorem 1.9, as follows:
Theorem 4.6.
Let $r,k\in\mathbb{Z}_{>0}$ with $r\geq 2$ and $R(z)\in\mathbb{C}[z]\setminus\{0\}$. Set
$P(z)=z^{r}R(z^{k})\quad\text{and}\quad Q(z)=z^{r}R(z)^{k}.$
Assume that $\gcd(r(r^{M}-1),k)=1$ for some $M\in\mathbb{Z}_{>0}$. Then there exist integers $d>c_{1}>0$ such that for every $N\in\mathbb{Z}_{\geq 0}$,
$S_{c_{1}+Nd}(P)=S_{c_{1}+Nd}(Q).$
Remark 4.7.
According to the proof of Theorem 4.6, we can take $c_{1}$ to be the minimal positive integer such that $\gcd(r(r^{c_{1}}-1),k)=1$ and $d$ to be the minimal positive integer with $r^{d}\equiv 1\pmod{k}$.
Proof.
If $k=1$, then $P=Q$ and the conclusion is trivial. Assume $k\geq 2$. Write $R(z)=z^{l}\cdot R_{0}(z)$, where $l\in\mathbb{Z}_{\geq 0}$ and $R_{0}(z)\in\mathbb{C}[z]$ with $R_{0}(0)\neq 0$. For $r_{0}=r+lk$ and $M\in\mathbb{Z}_{>0}$ such that $\gcd(r(r^{M}-1),k)=1$, we also have $\gcd(k,r_{0})=\gcd(r_{0}^{M}-1,k)=1$. Replacing $r$ and $R$ by $r_{0}$ and $R_{0}$, respectively, we may assume $R(0)\neq 0$. If $\deg(R)=0$, then both $P$ and $Q$ are conjugate to $z^{r}$ (recall $r\geq 2$), and the conclusion is immediate. Henceforth assume $\deg(R)=s\geq 1$.
First, for every polynomial $f(z)\in\mathbb{C}[z]$ of degree $m\geq 2$ and every $n\in\mathbb{Z}_{>0}$, the point $\infty$ has multiplicity $1$ in ${\rm Fix}(f^{\circ n})$ with $\rho_{f^{\circ n}}(\infty)=0$. Thus we only need to consider periodic points in $\mathbb{C}$. Let ${\rm Fix}(\cdot)$ denote ${\rm Fix}(\cdot,\mathbb{C})$ from now on.
Fix an arbitrary integer $n\geq 1$. From the commutative relation
(4.9) $Q^{\circ n}\circ z^{k}=z^{k}\circ P^{\circ n},$
it is easy to see that the map (between multi-sets)
$h:{\rm Fix}(P^{\circ n})\to{\rm Fix}(Q^{\circ n}),\ z_{0}\mapsto z_{0}^{k}$
is well-defined.
Set $F_{1}={\rm Fix}(P^{\circ n})$ and $F_{2}={\rm Fix}(Q^{\circ n})$. For a multi-set $X$ and an element $x\in X$, let ${\rm multi}_{X}(x)$ denote the multiplicity of $x$ in $X$. We claim that for every $z_{0}\in F_{1}$, we have
(4.10) ${\rm multi}_{F_{1}}(z_{0})={\rm multi}_{F_{2}}(z_{0}^{k})\quad\text{and}\quad\rho_{P^{\circ n}}(z_{0})=\rho_{Q^{\circ n}}(z_{0}^{k}).$
Recall that for every rational map $f\in\mathbb{C}(z)$ of degree at least $2$ and a fixed point $z_{0}\in{\rm Fix}(f)$, we have $\rho_{f}(z_{0})\neq 1$ if and only if ${\rm multi}_{{\rm Fix}(f)}(z_{0})=1$. Differentiating (4.9) and evaluating at $z_{0}\in{\rm Fix}(P^{\circ n})$, we obtain
(4.11) $\left((Q^{\circ n})^{\prime}(z_{0}^{k})-(P^{\circ n})^{\prime}(z_{0})\right)\cdot z_{0}^{k-1}=0.$
Let $z_{0}\in{\rm Fix}(P^{\circ n})\setminus\{0\}$. By (4.11),
$\rho_{P^{\circ n}}(z_{0})=(P^{\circ n})^{\prime}(z_{0})=(Q^{\circ n})^{\prime}(z_{0}^{k})=\rho_{Q^{\circ n}}(z_{0}^{k}).$
Thus (4.10) holds when $\rho_{P^{\circ n}}(z_{0})=\rho_{Q^{\circ n}}(z_{0}^{k})\neq 1$. Assume $\rho_{P^{\circ n}}(z_{0})=1$. Writing $\zeta_{k}=\exp(2\pi i/k)$, we have
(4.12) $Q^{\circ n}(z^{k})-z^{k}=(P^{\circ n}(z))^{k}-z^{k}=(P^{\circ n}(z)-z)\prod_{j=1}^{k-1}(P^{\circ n}(z)-\zeta_{k}^{j}z).$
The left-hand side of (4.12) can be regarded as a polynomial in the variable $z^{k}-z_{0}^{k}$, and the right-hand side as a polynomial in $z-z_{0}$. Since $z_{0}\neq 0$, we have ${\rm ord}_{z-z_{0}}(z^{k}-z_{0}^{k})=1$. Applying $\rm ord$} to both sides of (4.12), we obtain
${\rm ord}_{z-z_{0}^{k}}(Q^{\circ n}(z)-z)$ $={\rm ord}_{z^{k}-z_{0}^{k}}(Q^{\circ n}(z^{k})-z^{k})\cdot 1$
$={\rm ord}_{z^{k}-z_{0}^{k}}(Q^{\circ n}(z^{k})-z^{k})\cdot{\rm ord}_{z-z_{0}}(z^{k}-z_{0}^{k})$
$={\rm ord}_{z-z_{0}}(Q^{\circ n}(z^{k})-z^{k})$
$={\rm ord}_{z-z_{0}}(P^{\circ n}(z)-z).$
Thus ${\rm multi}_{F_{1}}(z_{0})={\rm multi}_{F_{2}}(z_{0}^{k})\geq 2$.
For $z_{0}=0=z_{0}^{k}$, since $r\geq 2$, 0 is a super-attracting fixed point of both $P^{\circ n}$ and $Q^{\circ n}$, i.e., $\rho_{P^{\circ n}}(0)=\rho_{Q^{\circ n}}(0)=0$. Hence ${\rm multi}_{F_{1}}(0)={\rm multi}_{F_{2}}(0)=1$, concluding the proof of (4.10).
From (4.10), if $h(z_{0})=h(z_{1})$ implies $z_{0}=z_{1}$ for $z_{0},z_{1}\in F_{1}$ (in which case we say $n$ is “good”), then $S_{n}(P)=S_{n}(Q)$. We show that $n$ is good for all $n$ in an infinite arithmetic progression.
Assume $n\in\mathbb{Z}_{>0}$ is not good. Then there exist $z_{0}\neq z_{1}$ in $F_{1}$ with $z_{1}^{k}=z_{0}^{k}$. Hence $z_{1}\neq 0$ and $z_{1}=\zeta z_{0}$ for some $\zeta\in\mathbb{C}$ with $\zeta^{k}=1$ and $\zeta\neq 1$. It is easy to see that there exists $R_{n}(z)\in\mathbb{C}[z]$ with $R_{n}(0)\neq 0$ such that $P^{\circ n}(z)=z^{r^{n}}\cdot R_{n}(z^{k})$ and $Q^{\circ n}(z)=z^{r^{n}}\cdot R_{n}(z)^{k}$. From $P^{\circ n}(z_{0})=z_{0}$ and $P^{\circ n}(\zeta z_{0})=\zeta z_{0}$, we deduce $\zeta^{r^{n}}=\zeta$, so
$1<{\rm ord}_{\mathbb{C}^{\times}}(\zeta)\mid\gcd(r^{n}-1,k).$
Thus, for every integer $n\geq 1$, either $n$ is good, or $\gcd(r^{n}-1,k)>1$.
Let $M$ be the minimal positive integer such that $\gcd(r(r^{M}-1),k)=1$. In particular, $r$ and $k$ are coprime. By Euler’s theorem, $r^{\varphi(k)}\equiv 1\pmod{k}$. Let $d$ be the minimal positive integer with $r^{d}\equiv 1\pmod{k}$. Then $d\mid\varphi(k)$. For every integer $N\geq 0$, we have
$r^{M+Nd}-1\equiv r^{M}-1\pmod{k}$,
so $\gcd(r^{M+Nd}-1,k)=1$. Therefore, every $n$ in the arithmetic progression $(M+Nd)_{N=0}^{\infty}$ is good. Since the sequence $(r^{n}\bmod k)_{n}$ has exact period $d$, $M\leq d$ by minimality. In fact, we must have $M<d$ (otherwise $k\mid\gcd(r^{d}-1,k)=\gcd(r^{M}-1,k)=1$, contradicting $k>1$). ∎
Example 4.8.
Given integers $n\geq 1$ and $k\geq 2$, let $V(z)\in\mathbb{C}[z]$ be a polynomial of degree $n$ with $V(0)\neq 0$. Consider the Ritt move
$(P(z):=z\cdot V(z^{k}),R(z):=z\cdot V(z)^{k}).$
We show that for general such $V$, we have $S_{1}(P)\neq S_{1}(R)$. Write $V(z)=a_{n}z^{n}+\cdots+a_{1}z+a_{0}$ with $(a_{n},\dots,a_{0})\in\mathbb{C}^{\times}\times\mathbb{C}^{n-1}\times\mathbb{C}^{\times}$. It is easy to see that $S_{1}(R)$ is the multi-set
$\{0,\rho_{R}(0)=a_{0}^{k},\rho_{R}(w_{1})=1+kw_{1}\frac{V^{\prime}(w_{1})}{V(w_{1})},\dots,\rho_{R}(w_{nk})=1+kw_{nk}\frac{V^{\prime}(w_{nk})}{V(w_{nk})}\},$
where $V(z)^{k}=a(z-w_{1})\cdots(z-w_{nk})+1$. One can check that
$\prod_{j=1}^{nk}\left(a_{0}-1-k\,w_{j}\frac{V^{\prime}(w_{j})}{V(w_{j})}\right)$
is a nonzero polynomial in $a_{0},\dots,a_{n}$; denote it by $f(a_{0},\dots,a_{n})$. When $(a_{n},\dots,a_{0})$ does not lie on the union of the hypersurfaces $a_{0}=a_{0}^{k}$ and $f(a_{0},\dots,a_{n})=0$, we have $\rho_{P}(0)=a_{0}\notin S_{1}(R)$, hence $S_{1}(P)\neq S_{1}(R)$. Therefore, for $(a_{n},\dots,a_{0})$ in a Zariski-dense open subset of $\mathbb{C}^{\times}\times\mathbb{C}^{n-1}\times\mathbb{C}^{\times}$, we have $S_{1}(P)\neq S_{1}(R)$.
Example 4.9.
Let $P(z)=z^{2}(z^{3}+1)$ and $Q(z)=z^{2}(z+1)^{3}$, which satisfy the hypothesis of Theorem 4.6. Indeed, by Remark 4.7, for every odd positive integer 2n12n-1 ($n\in\mathbb{Z}_{>0}$), we have
$S_{2n-1}(P)=S_{2n-1}(Q).$
Using computer computations, one finds that $S_{2}(P)\neq S_{2}(Q)$ and $S_{4}(P)\neq S_{4}(Q)$, so $P$ and $Q$ do not have the same multiplier spectrum.
Example 4.10.
Consider the Ritt move $(P(z)=z(z^{2}-3),R(z)=z(z-3)^{2})$. Then P(z2)+2=R(z)P(z-2)+2=R(z), hence $P$ and $R$ are conjugate and have the same multiplier spectrum.
Beyond the coincidence of the multiplier spectra on an infinite arithmetic progression under additional hypotheses, we can say more about the relation of the multiplier spectra in case (2) of Theorem 1.9.
Proposition 4.11.
Let $r,k\in\mathbb{Z}_{>0}$ with $r\geq 2$ and $R(z)\in\mathbb{C}[z]\setminus\{0\}$. Set
$P(z)=z^{r}\cdot R(z^{k})\quad\text{and}\quad Q(z)=z^{r}\cdot R(z)^{k}.$
  1. (1)
    For every $z_{0}\in{\rm Per}(P)$, there exist $w_{0}\in{\rm Per}(Q)$ and $m\in\mathbb{Z}_{>0}$ such that $\rho_{P}(z_{0})=\rho_{Q}(w_{0})^{m}$ and $n_{Q}(w_{0})\mid n_{P}(z_{0})$.
  2. (2)
    For every $w_{0}\in{\rm Per}(Q)$, there exist $z_{0}\in{\rm Per}(P)$ and $m\in\mathbb{Z}_{>0}$ such that $\rho_{Q}(w_{0})=\rho_{P}(z_{0})^{m}$ and $n_{Q}(w_{0})\mid n_{P}(z_{0})$.
In fact, in both (1) and (2) we can take $m=n_{P}(z_{0})/n_{Q}(w_{0})$.
Proof.
  1. (1)
    Let $z_{0}\in{\rm Per}(P)$ with exact period $l$. If $\rho_{P}(z_{0})=0$, take $w_{0}=\infty$ and $m=1$. Assume $\rho_{P}(z_{0})\neq 0$. In particular, $z_{0}\notin\{0,\infty\}$ since $r\geq 2$. From (4.11) (which holds without assuming $\gcd(r(r^{M}-1),k)=1$), we have n_Q(z_0^k)∣l and ρ_Q^∘l(z_0^k)=ρ_P^∘l(z_0)=ρ_P(z_0). Set $w_{0}=z_{0}^{k}$ and $m=l/n_{Q}(w_{0})$.
  2. (2)
    Let $w_{0}\in{\rm Per}(Q)$ with exact period $l$. If $\rho_{Q}(w_{0})=0$, take $z_{0}=\infty$ and $n=m=1$. Assume that $\rho_{Q}(w_{0})\neq 0$. In particular, $w_{0}\notin\{0,\infty\}$. By [FG22, Lemma 3.42], there exists $z_{0}\in{\rm PrePer}(P)$ such that $z_{0}^{k}=w_{0}$. Replacing $z_{0}$ by $P^{\circ l}(z_{0})$ for some suitable $l\geq 1$ if necessary, we may assume $z_{0}\in{\rm Per}(P)$. From (4.11), the choices $z_{0}$ and $m=n_{P}(z_{0})/l$ satisfy the requirement.
Remark 4.12.
We can give more explicit descriptions of $m$ and relations between $(S_{n}(P))$ and $(S_{n}(Q))$. For example, in the proof of 2, we must have
$P^{\circ jn}(z_{0})=\zeta_{j}z_{0}\quad(j=0,\dots,m-1),$
where $1=\zeta_{0},\dots,\zeta_{m-1}$ are distinct $k$-th roots of unity. In particular, $m\leq k$. We can deduce some other restrictions involving $\zeta_{j}$. Similarly, we have $m\leq k$ in 1.
Proposition 4.11 shows that in the Ritt move case, the multipliers of $P$ and $Q$ are mutually multiplicatively dependent. It is natural to consider multipliers modulo multiplicative dependence. For every rational map $f\in\mathbb{C}(z)$ of degree at least $2$, let $V(f)$ denote the $\mathbb{Q}$-vector space generated by $\{\chi_{f}(z):z\in{\rm Per}^{*}(f)\}\subseteq\mathbb{R}$, which is studied by the author, Ji, and Xie in [JXZ25].
Corollary 4.13.
Let $r,k\in\mathbb{Z}_{>0}$ with $r\geq 2$ and $R(z)\in\mathbb{C}[z]\setminus\{0\}$. Set
$P(z)=z^{r}\cdot R(z^{k})\quad\text{and}\quad Q(z)=z^{r}\cdot R(z)^{k}.$
Then the following hold:
  1. (1)
    $V(P)=V(Q)$.
  2. (2)
    $P$ is exceptional if and only if $Q$ is exceptional.
  3. (3)
    $P$ is PCF if and only if $Q$ is PCF.
Proof.
(1) This follows from Proposition 4.11 and the definition of $V(\cdot)$.
(2) This follows from (1) and [JXZ25, Theorem 1.5]. (The converse of [JXZ25, Theorem 1.5] holds; see [Mil06].)
(3) This follows from Proposition 4.11 and [JXZ25, Theorem 1.14]. ∎
Remark 4.14.
Part (2) of Corollary 4.13 is a direct corollary of [Pak17, Theorem 4.4] (see also [FG22, Theorem 3.39]), which states: for two polynomials $f,g\in\mathbb{C}[z]$ of degree $d\geq 2$ with $\geq g, ff$ is conjugate to $z^{d}$ if and only if $g$ is conjugate to $z^{d}$, and $f$ is conjugate to $\pm T_{d}$ if and only if $g$ is conjugate to $\pm T_{d}$. Here $T_{d}(z)$ is the Chebyshev polynomial of degree $d$, and the exceptional polynomials of degree $d$ are exactly $z^{d},\pm T_{d}(z)$ up to conjugacy.
Remark 4.15.
Let $f,g\in\mathbb{C}(z)$ be two rational maps of the same degree $d\geq 2$. Assume first that $f\geq g$. It is easy to see that $V(f)\cap V(g)+W=V(f)+V(g)$ in $\mathbb{R}$ for some finite-dimensional subspace $W$ of $V(f)+V(g)$. Consequently, for intertwined $f$ and $g$, we have $V(f)\cap V(g)+W=V(f)+V(g)$ in $\mathbb{R}$ for some finite-dimensional subspace $W$ of $V(f)+V(g)$. Combined with [JXZ25, Theorem 1.3], we obtain that $f$ is exceptional if and only if $g$ is.
5. Further directions and problems
5.1. Multiplier spectrum over an arithmetic progression
We can ask to what extent the converse of Theorem 4.6 holds.
Question 5.1.
Let $P(z),Q(z)\in\mathbb{C}[z]$ be (non-exceptional) polynomials of the same degree $M\geq 2$. Under what conditions do there exist positive integers $c_{1}<d$ such that
$S_{c_{1}+Nd}(P)=S_{c_{1}+Nd}(Q)$
for every $N\in\mathbb{Z}_{\geq 0}$? Are there examples that do not arise from Theorem 4.6, up to equivalence?
Note that in Theorem 4.6 and Question 5.1, the first term $c_{1}$ is strictly less than the common difference $d$ of the arithmetic progression. When $c_{1}\geq d$, we can always find $k\in\mathbb{Z}_{>0}$ such that $c_{1}<dk$ and then consider the arithmetic progression $(c_{1}+Ndk)_{N=0}^{\infty}$ instead. It is interesting to consider when we can take $c_{1}=d$, which relates to the following notion of stable multiplier spectrum:
Definition 5.2 (Stable multiplier spectrum).
Let $f,g\in\mathbb{C}(z)$ be two rational maps of the same degree $d\geq 2$. W say $f$ and $g$ have the same stable multiplier spectrum if $\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})$ for some $k\in\mathbb{Z}_{>0}$.
Remark 5.3.
Clearly, $f$ and $g$ have the same stable multiplier spectrum if and only if there exists an arithmetic progression $A=(Nk)_{N=1}^{\infty}$ with the first term and common difference both equal to some integer $k\in\mathbb{Z}_{>0}$ such that $S_{n}(f)=S_{n}(g)$ for every $n\in A$. Such an arithmetic progression $A=(Nk)_{N=1}^{\infty}\subseteq\mathbb{Z}_{>0}$ is called equidistant.
Example 5.4.
If there exists $k\in\mathbb{Z}_{>0}$ such that $f^{\circ k}=g^{\circ k}$, then $f$ and $g$ have the same stable multiplier spectrum. For example, for $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$ and $n\in\mathbb{Z}_{>0}$, the polynomials $P(z)=zV(z^{2n})$ and $Q(z)=-zV(z^{2n})$ (which are a specific case of Theorem 1.9 $(2)$) have the same stable multiplier spectrum, since $P^{\circ 2}=Q^{\circ 2}$.
Fix an integer $d\geq 2$. Let
$E_{d}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C}):\exists k\in\mathbb{Z}_{>0},\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})\}.$
Then $E_{d}=\cup_{k\geq 1}E_{d,k}$ is a countable union of Zariski closed sets, where
$E_{d,k}=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C}):\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})\}$
is Zariski closed. (Note that the map
$\kappa:\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathcal{M}_{d^{k}}(\mathbb{C})\times\mathcal{M}_{d^{k}}(\mathbb{C}),([f],[g])\mapsto([f^{\circ k}],[g^{\circ k}])$
is an algebraic morphism and $E_{d,k}$ is the inverse image of $R_{d^{k},m_{d^{k}}}$ (see §1.1) under $\kappa$.) Note that for positive integers $k_{1}$ and $k_{2}$ with $k_{1}\mid k_{2}$, we have $E_{d,k_{1}}\subseteq E_{d,k_{2}}$. Hence $E_{d}$ can be written as a countable union of increasing Zariski closed subsets of $\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})$. Currently, it is unclear whether $E_{d}$ is a Zariski closed subset.
Question 5.5.
For $d\geq 2$, is $E_{d}$ Zariski closed in $\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})$?
Remark 5.6.
The notion of stable multiplier spectrum and the Question 5.5 are proposed by Prof. Xie and we learn them from private communications with Prof. Xie.
Replacing “multiplier” by “length” in the above discussion yields the analogous notion of stable length spectrum.
We show that the multiplier spectrum morphism over an equidistant arithmetic progression remains generically injective on the moduli space.
Proposition 5.7.
Let $d\geq 2$ and $n\geq 1$ be integers. The morphism
$\tau_{d}[n]:\mathcal{M}_{d}(\mathbb{C})\to\prod_{j=1}^{m_{d^{n}}}\mathbb{A}^{N_{d^{n},j}}(\mathbb{C}),[f]\mapsto\tau_{d^{n}}([f^{\circ n}])$
is generically injective. Similarly, the restriction $\tilde{\tau}_{d}[n]$ of $\tau_{d}[n]$ to ${\rm MPoly}^{d}(\mathbb{C})$ is also generically injective.
Proof.
Let $\kappa_{d,n}:\mathcal{M}_{d}\to\mathcal{M}_{d^{n}}$ be the composition morphism $f]\mapsto[f^{\circ n}]$. Then $\tau_{d}[n]=\tau_{d^{n}}\circ\kappa_{d,n}$. By Theorem 1.3, $\tau_{d^{n}}$ is generically injective. Thus, it suffices to show $\kappa_{d,n}$ is also generically injective. This follows from [Pak25b, Theorem 5.4].
To show the generic injectivity of $\tilde{\tau}_{d}[n]$, replacing Theorem 1.3 by Theorem 1.4, it suffices to show that the composition morphism
$\tilde{\kappa}_{d,n}:{\rm MPoly}^{d}(\mathbb{C})\to{\rm MPoly}^{d^{n}}(\mathbb{C}),[f]\mapsto[f^{\circ n}]$
is also generically injective. If $d=2$, then Theorem 3.25 (2) implies the generic injectivity of $\tilde{\kappa}_{2,n}$. Assume now $d\geq 3$. By Lemmas 3.15 and 3.16, there exists a nonempty Zariski open subset $U_{d}\subseteq{\rm MPoly}^{d}(\mathbb{C})$ such that for every $\in U_{d}, ff$ is pre-simple and $G(f)=1$. We claim that $\tilde{\kappa}_{d,n}|_{U_{d}}$ is injective. Let $f],[g]\in U_{d}$ such that $\tilde{\kappa}_{d,n}([f])=\tilde{\kappa}_{d,n}([g])$. After replacing $g$ by a suitable affine conjugate if necessary, we may assume $f^{\circ n}=g^{\circ n}$. Since $f$ is pre-simple of degree $d\geq 3$, Corollary 3.9 implies that $\tau\circ f=g=f\circ\sigma$ for some $\sigma,\tau\in\mathbb{C}[z]$ of degree one. Since $G(f)=1$, we obtain $\sigma=1$, hence $f=g$. Thus, the morphism $\tilde{\kappa}_{d,n}$ is injective on $U_{d}$. Therefore, $\tilde{\tau}_{d}[n]$ is generically injective. ∎
The above theorem immediately implies the following corollary:
Corollary 5.8.
Let $d\geq 4$ be an integer. The stable multiplier map on $\mathcal{M}_{d}(\mathbb{C})$ is very generically injective; that is, there exists a subset $V\subseteq\mathcal{M}_{d}(\mathbb{C})$ of the form
$V=\mathcal{M}_{d}(\mathbb{C})\setminus\left(\cup_{m=1}^{\infty}Z_{m}\right),$
where each $Z_{m}$ is a proper Zariski closed subset of $\mathcal{M}_{d}(\mathbb{C})$, such that for all $f],[g]\in V$, if $f$ and $g$ have the same stable multiplier spectrum, then $[f]=[g]$.
We may ask whether the multiplier spectrum over any infinite arithmetic progression is generically injective on the moduli space for further study. (Note that generalized Lattès maps are sparse in the moduli space, and hence so are maps arising from the Ritt move process.)
5.2. Non-injective locus and length spectrum
We aim to provide a more precise description of the non-injective locus than that given in Theorem 1.9, in light of Conjecture 1.3.
There are many directions for possible generalizations of Theorem 1.9. We list some of them as follows:
  • Can we bound $N$ and #S\#S in Theorem 1.9 explicitly?
  • Can we exclude case (2) in light of Conjecture 1.3?
  • Can we describe ${\rm PNI}_{d}$ more precisely (possibly ignoring zero-dimensional components)?
  • What can be said for rational maps (not just polynomials) with the same multiplier spectrum?
There are less known results for the length spectrum than the multiplier spectrum, since the length spectrum contains less information and the failure of being an (algebraic) morphism between schemes for the length spectrum map. It is more difficult to study the length spectrum. A significant result proved by Ji and Xie proved in this direction is that aside from the flexible Lattès family, the length spectrum determines the conjugacy class of rational maps up to finitely many choices [JX23, Theorem 1.5]. We can ask some further questions for the length spectrum as follows:
  • Does there exist a Zariski-dense open subset $U$ of $\mathcal{M}_{d}(\mathbb{C})$ such that every pair $(x,y)\in U\times U$ with $L(x)=L(y)$ must have the form $([f],[f])$ or $([f],[\overline{f}])$? Here $\overline{f}$ represents the rational map obtained by applying complex conjugate to all the coefficients of $f$. Note that a positive answer to the above question is equivalent to the correctness of a conjecture of Ji-Xie [JX25, Conjecture 1.9]. We can ask the same question with the weaker requirement that $U$ has full (or just positive) Lebesgue measure (which may not be Zariski-dense open). Similar questions can be asked for polynomials as well.
  • Classify rational maps (or polynomials) with the same length spectrum.